### v2

# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

Abstract
Recent studies have highlighted the benefits of generating multiple synthetic datasets for supervised learning, from increased accuracy to more effective model selection and uncertainty estimation. These benefits have clear empirical support, but the theoretical understanding of them is currently very light. We seek to increase the theoretical understanding by deriving bias-variance decompositions for several settings of using multiple synthetic datasets, including differentially private synthetic data. Our theory predicts multiple synthetic datasets to be especially beneficial for high-variance downstream predictors, and yields a simple rule of thumb to select the appropriate number of synthetic datasets in the case of mean-squared error and Brier score. We investigate how our theory works in practice by evaluating the performance of an ensemble over many synthetic datasets for several real datasets and downstream predictors. The results follow our theory, showing that our insights are practically relevant.
1 Introduction
Synthetic data has recently attracted significant attention for several applications in machine learning. The idea is to generate a dataset that preserves the population-level attributes of the real data. This makes the synthetic data useful for analysis, while also accomplishing a secondary task, such as improving model evaluation (van Breugel et al., 2023b), fairness (van Breugel et al., 2021), data augmentation (Antoniou et al., 2018; Das et al., 2022) or privacy (Liew et al., 1985; Rubin, 1993). For privacy, differential privacy (DP) (Dwork et al., 2006) is often combined with synthetic data generation (Hardt et al., 2012; McKenna et al., 2021) to achieve provable privacy protection, since releasing synthetic data without DP can be vulnerable to membership inference attacks (Stadler et al., 2022; van Breugel et al., 2023c; Meeus et al., 2023).
Several lines of work have considered generating multiple synthetic datasets from one real dataset for various purposes, including statistical inference (Rubin, 1993; Raghunathan et al., 2003; Räisä et al., 2023a) and supervised learning (van Breugel et al., 2023a), the latter of which is our focus.
In supervised learning, van Breugel et al. (2023a) proposed using multiple synthetic datasets in the form of an ensemble. They propose training a predictive model separately on each synthetic dataset, and ensembling these models by averaging their predictions, which they call a deep generative ensemble (DGE). We drop the word “deep” in this paper and use the term generative ensemble, as we do not require the generator to be deep in any sense. The DGE was empirically demonstrated to be beneficial in several ways by van Breugel et al. (2023a), including predictive accuracy, model evaluation, model selection, and uncertainty estimation. Followup work has applied DGEs to improve model evaluation under distribution shifts and for small subgroups (van Breugel et al., 2023b).
However, van Breugel et al. (2023a) have very little theoretical analysis of the generative ensemble. Their theoretical justification assumes the data is generated from a posterior predictive distribution. This assumption can be justified heuristically for deep generative models through a Bayesian interpretation of deep ensembles (Lakshminarayanan et al., 2017; Wilson & Izmailov, 2020; Wilson, 2021). However, this justification only applies to generators with highly multi-modal losses, like neural networks. It also does not provide any insight on how different choices in the setup, like the choice of downstream predictor, affect the performance of the ensemble.
The bias-variance decomposition (Geman et al., 1992) and its generalisations (Ueda & Nakano, 1996; Wood et al., 2023) are classical tools that provide insight into supervised learning problems. The standard bias-variance decomposition from Geman et al. (1992) considers predicting $y\in\mathbb{R}$ given features $x\in\mathcal{X}$, using predictor $g(x;D)$ that receives training data $D$. The mean-squared error (MSE) of $g$ can be decomposed into bias, variance, and noise terms:
$\underbrace{\operatorname*{\mathbb{E}}_{D,y}[(y-g(x;D))^{2}]}_{\mathrm{MSE}}=\underbrace{(f(x)-\operatorname*{\mathbb{E}}_{D}[g(x;D)])^{2}}_{\mathrm{Bias}}+\underbrace{\operatorname*{Var}_{D}[g(x;D)]}_{\mathrm{Variance}}+\underbrace{\operatorname*{Var}_{y}[y]}_{\mathrm{Noise}}$ (1)
where we $f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor. $x$ is considered fixed, so all the random quantities in the decomposition are implicitly conditioned on $x$. While MSE is typically only used with regression problems, the decomposition also applies to the Brier score (Brier, 1950) in classification, which is simply the MSE of class probabilities.
We seek to provide deeper theoretical understanding of generative ensembles through bias-variance decompositions, which provide a more fine-grained view of how different choices in the setup affect the end result.
Contributions
  1. 1.
    We derive a bias-variance decomposition for the MSE or Brier score of generative ensembles under an i.i.d. assumption in Theorem 2.2. This decomposition is simply a sum of interpretable terms, which makes it possible to predict how various choices affect the error. In particular, the number of synthetic datasets only affects the variance-related terms in the decomposition, predicting that generative ensembles work best with high-variance and low bias models, like interpolating decision trees.
  2. 2.
    We derive several practical considerations from our decomposition, including a simple rule of thumb to select the number of synthetic datasets in Section 2.3. In summary, 2 synthetic datasets give 50% of the potential benefit from multiple synthetic datasets, 10 give 90% of the benefit and 100 give 99% of the benefit. This also applies to bagging ensembles like random forests, which is likely to be of independent interest.
  3. 3.
    We generalise the decomposition of Theorem 2.2 to differentially private (DP) generation algorithms that do not split their privacy budget between the multiple synthetic datasets111 Theorem 2.2 applies to DP algorithms that split the privacy budget, but it is not clear if multiple synthetic datasets are beneficial with these algorithms, as splitting the privacy budget between more synthetic datasets means that each one requires adding more noise, degrading the quality of the synthetic data. in Theorem 2.4, to non-i.i.d. synthetic data in Appendix B, and to Bregman divergences in Appendix C.
  4. 4.
    We evaluate the performance of a generative ensemble on several datasets, downstream prediction algorithms, and error metrics in Section 3. The results show that out theory applies in practice: multiple synthetic datasets generally decrease all of the error metrics, especially for high-variance models where the theory predicts multiple synthetic datasets to have the highest impact. In addition, our rule of thumb makes accurate predictions when the error can be accurately estimated.
1.1 Related Work
Ensembling generative models has been independently proposed several times (Wang et al., 2016; Choi et al., 2019; Luzi et al., 2020; Chen et al., 2023; van Breugel et al., 2023a). The inspiration of our work comes from van Breugel et al. (2023a), who proposed ensembling predictive models over multiple synthetic datasets, and empirically studied how this improves several aspects of performance in classification.
Generating multiple synthetic datasets has also been proposed with statistical inference in mind, for both frequentist (Rubin, 1993; Raghunathan et al., 2003; Räisä et al., 2023b), and recently Bayesian (Räisä et al., 2023a) inference. These works use the multiple synthetic datasets to correct statistical inferences for the extra uncertainty from synthetic data generation.
In a similar manner, multiple imputation (Rubin, 1987; Murray, 2018) imputes missing values multiple times to a dataset, which allows correcting uncertainty estimates for the extra uncertainty from missing values. Multiple imputation is typically used with statistical inference in mind, but the imputed datasets could be used to train an ensemble.
The bias-variance decomposition was originally derived by Geman et al. (1992) for standard regressors using MSE as the loss. James (2003) generalised the decomposition to symmetric losses, and Pfau (2013) generalised it to Bregman divergences.
Ueda & Nakano (1996) were the first to study the MSE bias-variance decomposition with ensembles, and Gupta et al. (2022); Wood et al. (2023) later extended the ensemble decomposition to other losses. All of these also apply to generative ensembles, but they only provide limited insight for them, as they do not separate the synthetic data generation-related terms from the downstream-related terms.
2 Bias-Variance Decompositions for Generative Ensembles
In this section, we make our main theoretical contributions. We start by defining our setting in Section 2.1. Then we derive a bias-variance decomposition for generative ensembles in Section 2.2 and derive a simple rule of thumb that can be used to select the number of synthetic datasets in Section 2.3. The generative ensemble decomposition does not apply to several differentially private synthetic data generation methods that generate all synthetic datasets conditional on a single privatised value, so we generalise the decomposition to apply to those in Section 2.4. We also present a third decomposition that applies to Bregman divergences, but is not as informative as the others, in Appendix C.
2.1 Problem Setting
We consider multiple synthetic datasets $D_{s}^{1:m}$, each of which is used to train an instance of a predictive model $g$. These models are combined into an ensemble $\hat{g}$ by averaging, so
$\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i}).$ (2)
We allow $g$ to be random, so $g$ can for example internally select among several predictive models to use randomly.
We assume that the synthetic datasets are generated by first i.i.d. sampling parameters $\theta_{1:m}|D_{r}$ given the real data, and then sampling synthetic datasets $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ given the parameters. This is how synthetic datasets are sampled in DGE (van Breugel et al., 2023a), where $\theta_{i}$ are the parameters of the generative neural network from independent training runs. This also encompasses bootstrapping, where $\theta_{i}$ would be the real dataset [^2] The $\theta_{i}$ random variables are i.i.d. if they are deterministically equal., and $p(D_{s}|\theta_{i})$ is the bootstrapping.
We will also consider a more general setting that applies to some differentially private synthetic data generators that do not fit into this setting in Section 2.4. In Appendix B, we consider settings without the i.i.d. assumption.
2.2 Mean-squared Error Decomposition
{theorem}
[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta|D_{r})$, $i=1,\dots,m$, be i.i.d. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then the mean-squared error in predicting $y$ from $x$ decomposes into six terms: model variance (MV), synthetic data variance (SDV), real data variance (RDV), synthetic data bias (SDB), model bias (MB), and noise $\operatorname*{Var}_{y}[y]$:
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],$ (3)
where
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\end{split}\quad\quad\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].\end{split}$ (4)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the distribution of the generator parameters, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$. The proofs of all theorems are in Appendix A.
We can derive the following practical considerations from Theorem 2.2:
  1. 1.
    There is a simple rule-of-thumb on how many synthetic datasets are beneficial: $m$ synthetic datasets give a $1-\frac{1}{m}$ fraction of the possible benefit from multiple synthetic datasets. For example, $m=2$ gives 50% of the benefit, $m=10$ gives 90% and $m=100$ gives 99%. This rule-of-thumb can be used to predict the MSE with many synthetic datasets from the results on just two synthetic datasets. More details in Section 2.3.
  2. 2.
    Increasing the number of synthetic datasets $m$ reduces the impact of MV and SDV. This means that high-variance models, like interpolating decision trees or 1-nearest neighbours, benefit the most from multiple synthetic datasets.
  3. 3.
    If $\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ is invariant to the size of the synthetic dataset, the only term that can change with the size of the synthetic dataset is MV. This means that the possible benefit of generating a single large synthetic dataset is smaller than generating many synthetic datasets, which can also reduce the impact of SDV.
  4. 4.
    The quality metrics for a synthetic data generator should include metrics that compare the distribution of the synthetic data to the real distribution, instead of just comparing a synthetic dataset to the real dataset. An example of the former is the downstream performance of predictive models, and examples of the latter are comparisons between the marginals of the real and synthetic datasets.
The last point can be derived by considering two extreme scenarios:
  1. 1.
    Generator is fitted perfectly, and generates from the real data generating distribution. Now
    $\begin{split}\mathrm{MV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{SDV}&=\mathrm{RDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (5)
    With one synthetic dataset, the result is the standard bias-variance trade-off. With multiple synthetic datasets, the impact of MV can be reduced, reducing the error compared to just using real data.
  2. 2.
    Return the real data: $D_{s}|D_{r}=D_{r}$ deterministically.333 Theorem 2.2 applies in this scenario even with multiple synthetic datasets, as the deterministically identical synthetic datasets are independent as random variables.
    $\begin{split}\mathrm{RDV}&=\operatorname*{Var}_{D_{r}}[g(x;D_{r})]\\ \mathrm{MV}&=\mathrm{SDV}=\mathrm{SDB}=0.\end{split}\quad\quad\begin{split}\mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-g(x;D_{r})]\\ \phantom{x}\end{split}$ (6)
    The result is the standard bias-variance trade-off. The number of synthetic datasets does not matter, as all of them would be the same anyway.
While both of these scenarios are unrealistic, they may be approximated by a well-performing algorithm. The generator from Scenario 2 is the optimal generator for metrics that compare the synthetic dataset to the real dataset, while the generator of Scenario 1 is optimal for metrics that compare the synthetic data and real data distributions. Multiple synthetic datasets are only beneficial in Scenario 1, which means that metrics comparing the distributions are more meaningful when multiple synthetic datasets are considered.
2.3 Estimating the effect of Multiple Synthetic Datasets
Next, we consider estimating the variance terms MV and SDV in Theorem 2.2 from a small number of synthetic datasets and a test set. These estimates can then be used to asses if more synthetic datasets should be generated, and how many more are useful.
We can simplify Theorem 2.2 to
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{Others},$ (7)
where Others does not depend on the number of synthetic datasets $m$. The usefulness of more synthetic datasets clearly depends on the magnitude of $\mathrm{MV}+\mathrm{SDV}$ compared to Others.
Since MSE depends on $m$, we can add a subscript to denote the $m$ in question: $\mathrm{MSE}_{m}$. Now (7) leads to the following corollary. {corollary} In the setting of Theorem 2.2, we have
$\mathrm{MSE}_{m}=\mathrm{MSE}_{1}-\left(1-\frac{1}{m}\right)(\mathrm{MV}+\mathrm{SDV}).$ (8)
Proof.
The claim follows by expanding $\mathrm{MSE}_{1}$ and $\mathrm{MSE}_{m}$ with (7). ∎
If we have two synthetic datasets, we can estimate $\mathrm{MV}+\mathrm{SDV}=2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, which gives the estimator
$\mathrm{MSE}_{m}=\mathrm{MSE}_{1}-2\left(1-\frac{1}{m}\right)(\mathrm{MSE}_{1}-\mathrm{MSE}_{2}).$ (9)
If we have more than two synthetic datasets, we can set $x_{m}=1-\frac{1}{m}$ and $y_{m}=\mathrm{MSE}_{m}$ in (8):
$y_{m}=y_{1}+x_{m}(\mathrm{MV}+\mathrm{SDV}),$ (10)
so we can estimate $\mathrm{MV}+\mathrm{SDV}$ from linear regression on $(x_{m},y_{m})$. However, this will likely have a limited effect on the accuracy of the $\mathrm{MSE}_{m}$ estimates, as it will not reduce the noise in estimating $\mathrm{MSE}_{1}$, which has a significant effect in (8).
From (8), we see that $\mathrm{MV}+\mathrm{SDV}$ is the maximum decrease in MSE that can be obtained from multiple synthetic datasets. This means that $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, or the linear regression estimate from (10), can be used as a diagnostic to check whether generating multiple synthetic datasets is worthwhile.
All terms in (7)-(10) depend on the target features $x$. We would like our estimates to be useful for typical $x$, so we will actually want to estimate $\operatorname*{\mathbb{E}}_{x}(\mathrm{MSE}_{m})$. Equations (7)-(10) remain valid if we take the expectation over $x$, so we can simply replace the MSE terms with their estimates that are computed from a test set.
Computing the estimates in practice will require that the privacy risk of publishing the test MSE is considered acceptable. The MSE for the estimate can also be computed from a separate validation set to avoid overfitting to the test set, but the risk of overfitting is small in this case, as $m$ has a monotonic effect on the MSE. Both of these caveats can be avoided by choosing $m$ using the rule of thumb that $m$ synthetic datasets give a $1-\frac{1}{m}$ of the potential benefit of multiple synthetic datasets, which is a consequence of (8).
Note that this MSE estimator can be applied to bagging ensembles (Breiman, 1996) like random forests (Breiman, 2001), since bootstrapping is a very simple form of synthetic data generation.
2.4 Differentially Private Synthetic Data Generators
Generating and releasing multiple synthetic datasets could increase the associated disclosure risk. One solution to this is differential privacy (DP) (Dwork et al., 2006; Dwork & Roth, 2014), which is a property of an algorithm that formally bounds the privacy leakage that can result from releasing the output of that algorithm. DP gives a quantitative upper bound on the privacy leakage, which is known as the privacy budget. Achieving DP requires adding extra noise to some point in the algorithm, lowering the utility of the result.
If the synthetic data is to be generated with DP, there are two possible ways to handle the required noise addition. The first is splitting the privacy budget across the $m$ synthetic datasets, and run the DP generation algorithm separately $m$ times. Theorem 2.2 applies in this setting. However, it is not clear if multiple synthetic datasets are beneficial in this case, as splitting the privacy budget requires adding more noise to each synthetic dataset. This also means that the rule of thumb from Section 2.3 will not apply. Most DP synthetic data generation algorithms would fall into this category (Aydore et al., 2021; Chen et al., 2020; Harder et al., 2021; Hardt et al., 2012; Liu et al., 2021; McKenna et al., 2019, 2021) if used to generate multiple synthetic datasets.
The second possibility is generating all synthetic datasets based on a single application of a DP mechanism. Specifically, a noisy summary $\tilde{s}$ of the real data is released under DP. The parameters $\theta_{1:m}$ are then sampled i.i.d. conditional on $\tilde{s}$, and the synthetic datasets are sampled conditionally on the $\theta_{1:m}$. This setting includes algorithms that release a posterior distribution under DP, and use the posterior to generate synthetic data, like the NAPSU-MQ algorithm (Räisä et al., 2023b) and DP variational inference (DPVI) (Jälkö et al., 2017, 2021).444 In DPVI, $\tilde{s}$ would be the variational approximation to the posterior.
The synthetic datasets are not i.i.d. given the real data in the second setting, so the setting described in Section 2.1 and assumed in Theorem 2.2 does not apply. However, the synthetic datasets are i.i.d. given the noisy summary $\tilde{s}$, so we obtain a similar decomposition as before. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta|\tilde{s})$, $i=1,\dotsc,m$, be i.i.d. given a DP summary $\tilde{s}$. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}|\theta_{i})$ independently, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})$. Then
$\begin{split}\mathrm{MSE}&=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}+\,(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],\end{split}$ (11)
where
$\begin{split}\mathrm{MSE}&=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]\\ \mathrm{MV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{SDV}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]\\ \mathrm{RDV}&=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})].\end{split}\quad\quad\begin{split}\mathrm{DPVAR}&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]\\ \mathrm{SDB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|\tilde{s}}[f_{\theta}(x)]]\\ \mathrm{MB}&=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]]\\ \phantom{x}\end{split}$ (12)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data, $\theta\sim p(\theta|D_{r})$ is a single sample from the distribution of the generator parameters, $D_{s}\sim p(D_{s}|\theta)$ is a single sample of the synthetic data generating process, and $f_{\theta}$ is the optimal predictor for the synthetic data generating process with parameters $\theta$. All random quantities are implicitly conditioned on $x$.
The takeaways from Theorem 2.4 are mostly the same as from Theorem 2.2, and the estimator from Section 2.3 also applies. The main difference is the DPVAR term in Theorem 2.4, which accounts for the added DP noise. As expected, the impact of DPVAR cannot be reduced with additional synthetic datasets.
3 Experiments
In this section, we describe our experiments. The common theme in all of them is generating synthetic data and evaluating the performance of several downstream prediction algorithms trained on the synthetic data. The performance evaluation uses a test set of real data, which is split from the whole dataset before generating synthetic data.
The downstream algorithms we consider are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), gradient boosted trees (GB), a multilayer perceptron (MLP) and a support vector machine (SVM) for both classification and regression. We also use linear regression (LR) and ridge regression (RR) on regression tasks, and logistic regression (LogR) on classification tasks, though we omit linear regression from the main text plots, as its results are nearly identical to ridge regression. Decision trees and 1-NN have a very high variance, as both interpolate the training data with the hyperparameters we use. Linear, ridge, and logistic regression have fairly small variance in contrast. Appendix D contains more details on the experimental setup, including details on the datasets, and downstream algorithm hyperparameters.
3.1 Effect of Multiple Synthetic Datasets
(a) Regression datasets
(b) Classification datasets
Figure 1: MSE on regression datasets (a) or Brier score on classification datasets (b) of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from synthpop. Increasing the number of synthetic datasets generally decreases both metrics, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), ridge regression (RR) and logistic regression (LogR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Tables S4 to S10 in the Appendix contain the numbers from the plots.
As our first experiment, we evaluate the performance of the synthetic data ensemble on 7 datasets, containing 4 regression and 3 classification tasks. See Appendix D.1 for details on the datasets. We use the synthetic data generators DDPM (Kotelnikov et al., 2023) and synthpop (Nowok et al., 2016), which we selected after a preliminary experiment described in Appendix E.1. We only plot the results from synthpop in the main text to save space, and defer the results of DDPM to Figure S2 in the Appendix. We generate 32 synthetic datasets, of which between 1 and 32 are used to train the ensemble. The results are averaged over 3 runs with different train-test splits.
On the regression datasets, our error metric is MSE, which is the subject of Theorem 2.2. The results in Figure 1(a) show that a larger number of synthetic datasets generally decreases MSE. The decrease is especially clear with downstream algorithms that have a high variance like decision trees and 1-NN. Low-variance algorithms like ridge regression have very little if any decrease from multiple synthetic datasets. This is consistent with Theorem 2.2, where the number of synthetic datasets only affects the variance-related terms.
On the classification datasets, we consider 4 error metrics. Brier score (Brier, 1950) is MSE of the class probability predictions, so Theorem 2.2 applies to it. Cross entropy is a Bregman divergence, so Theorem C.2 from Appendix C applies to it. We also included accuracy and area under the ROC curve (AUC) even though our theory does not apply to them, as they are common and interpretable error metrics, so it is interesting to see how multiple synthetic datasets affect them. We use their complements in the plots, so that lower is better for all plotted metrics. We only present the Brier score results in the main text in Figure 1(b), and defer the rest to Figures S3, S4 and S5 in Appendix E.2.
Because Theorem C.2 only applies to cross entropy when averaging log probabilities instead of probabilities, we compare both ways of averaging in the Appendix E.2, but only include probability averaging in the main text.
The results on the classification datasets in Figure 1(b) are similar to the regression experiment. A larger number of synthetic datasets generally decreases the score, especially for the high-variance models.
In Appendix E.4, we estimate the MV and SDV terms of the decomposition in Theorem 2.2. These confirm our claims on the model variances: decision trees and 1-NN have a high variance, while linear, ridge and logistic regression have a low variance.
(a) ACS 2018
(b) German credit
Figure 2: MSE or Brier score prediction on two datasets. The predictions are very accurate on these datasets. The solid lines for DDPM and synthpop (SP-P) show the same error MSE or Brier score as Figure 1, while the dashed lines show predicted MSE or Brier score. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. We omitted downstream algorithms with uninteresting flat curves from panel (a). See Figure S6 in the Appendix for the full panel (a), and Figures S7 and S8 for the other datasets. Tables S11 and S12 in the Appendix contain the numbers from the plots. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
3.2 Predicting performance from Two Synthetic Datasets
Next, we evaluate the predictions our rule of thumb from Section 2.3 makes. To recap, our rule of thumb predicts that the maximal benefit from multiple synthetic datasets is $2(\mathrm{MSE}_{1}-\mathrm{MSE}_{2})$, and $m$ synthetic datasets achieve a $1-\frac{1}{m}$ fraction of this benefit, as shown in (9).
To evaluate the predictions from the rule, we estimate the MSE on regression tasks and Brier score on classification tasks for one and two synthetic datasets from the test set. The setup is otherwise identical to Section 3.1, and the train-test splits are the same. We plot the predictions from the rule, and compare them with the measured test errors with more than two synthetic datasets.
Figure 2 contains the results for the ACS 2018 and German credit datasets, and Figures S6 to S8 in the Appendix contain the results for the other datasets. The predictions are very accurate on ACS 2018 and German credit, and reasonably accurate on the other datasets. The variance of the prediction depends heavily on the variance of the errors computed from the test data.
We also evaluated the rule on random forests without synthetic data, as it also applies to them. In this setting, the number of trees in the random forest is analogous to the number of synthetic datasets. We use the same datasets as in the previous experiments and use the same train-test splits of the real data. The results are in Figure S9 in the Appendix. The prediction is accurate when the test error is accurate, but can have high variance.
3.3 Differentially Private Synthetic Data
Figure 3: Brier score of the ensemble of downstream predictors with varying numbers of synthetic datasets $m$, generated with the DP methods AIM or NAPSU-MQ from the Adult dataset with a reduced set of features. Increasing the number of datasets generally decreases the score, even with AIM which splits the privacy budget between $m$ synthetic datasets. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are the same as in Figure 1. The black lines show the loss of the best non-DP downstream predictor trained on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Table S13 contains the numbers from the plots, and Figure S10 contains plots of the other error metrics.
In this experiment, we evaluate the performance of the generative ensemble in the setting of Theorem 2.4, where $m$ synthetic datasets are generated from a single noisy summary $\tilde{s}$ of the real data. We compare with splitting the privacy budget between $m$ synthetic datasets.
The algorithms we use are AIM (McKenna et al., 2022), which needs to split the privacy budget between the $m$ synthetic datasets, and NAPSU-MQ (Räisä et al., 2023b), which uses a single noisy summary. The dataset is the Adult dataset with a reduced set of columns, which is needed to keep the runtime of NAPSU-MQ reasonable. The downstream task is classification, so we use the same 4 metrics, and both probability and log-probability averaging, as in the non-DP classification experiment. We use fairly strict privacy parameters of $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. See Appendix D for the full details of the setup.
The results for Brier score are in Figure 3, and the results for all metrics are in Figure S10 in the Appendix. Increasing $m$ almost always improves the results, which is expected for NAPSU-MQ due to Theorem 2.4, but somewhat surprising for AIM, which splits the privacy budget. Once again, the high-variance predictors 1-NN, 5-NN and decision trees benefit the most from multiple synthetic datasets. AIM always outperforms NAPSU-MQ, which is likely a result of AIM being able to handle a more complicated set of input queries than NAPSU-MQ.
4 Discussion
Limitations
Our main theory assumes that the synthetic datasets are generated i.i.d. given either the real data or a DP summary. We generalise the bias-variance decompositions to non-i.i.d. settings in Appendix B. The implications of the decompositions, like the rule-of-thumb on the number of synthetic datasets, may not apply if i.i.d. assumption is removed.
Broader Impacts
Generating multiple synthetic datasets can potentially increase disclosure risk if appropriate precautions, such as differential privacy, are not used. Because of this, care should be taken and precautions like differential privacy should be used if possible when releasing multiple synthetic datasets of sensitive data, despite our theory suggesting that multiple synthetic datasets are beneficial for utility.
Conclusion
We derived bias-variance decompositions for using synthetic data in several cases: for MSE or Brier score with i.i.d. synthetic datasets given the real data (Section 2.2) and MSE with i.i.d. synthetic datasets given a DP summary of the real data (Section 2.4). We generalised these decompositions to non-i.i.d. synthetic datasets (Appendix B) and Bregman divergences (Appendix C). The decompositions make actionable predictions, such as yielding a simple rule of thumb that can be used to select the number of synthetic datasets. We empirically examined the performance of generative ensembles on several real datasets and downstream predictors, and found that the predictions of the theory generally hold in practice (Section 3). These findings significantly increase the theoretical understanding of generative ensembles, which is very limited in prior literature.
Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI as well as Grants 356499 and 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources.
Appendix A Missing Proofs
See 2.2
Proof.
With $m$ synthetic datasets $D_{s}^{1:m}$ and model $\hat{g}(x,D_{s}^{1:m})$ that combines the synthetic datasets, the classical bias-variance decomposition gives
$\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]=(f(x)-\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})])^{2}+\operatorname*{Var}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]+\operatorname*{Var}_{y}[y].$ (13)
Using the independence of the synthetic datasets, these can be decomposed further:
$\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g(x,D_{s}^{i})\right]=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x,D_{s})]=\operatorname*{\mathbb{E}}_{D_{r},D_{s}}[g(x,D_{s})],$ (14)
$\begin{split}\operatorname*{Var}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})\right]\\ &=\frac{1}{m^{2}}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}\left[g(x;D_{s})\right]\\ &=\frac{1}{m}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}|D_{r}}\left[g(x;D_{s})\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}\left[g(x;D_{s})\right],\end{split}$ (15)
and
$\operatorname*{Var}_{D_{s}|D_{r}}[g(x;D_{s})]=\operatorname*{\mathbb{E}}_{\theta|D_{r}}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]+\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x,D_{s})].$ (16)
The bias can be decomposed with $f_{\theta}(x)$:
$\begin{split}f(x)-\operatorname*{\mathbb{E}}_{D_{r},D_{s}}[g(x;D_{s})]&=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f(x)-f_{\theta}(x)+f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]+\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]]\end{split}$ (17)
Combining all of these gives the claim. ∎
See 2.4
Proof.
Using $\tilde{s}$ in place of the real data in Theorem 2.2 gives
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\operatorname*{Var}_{D_{r},\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],$ (18)
where
$\mathrm{MSE}$ $=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]$ (19)
$\mathrm{MV}$ $=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]$ (20)
$\mathrm{SDV}$ $=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ (21)
$\mathrm{SDB}$ $=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|\tilde{s}}[f_{\theta}(x)]]$ (22)
$\mathrm{MB}$ $=\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].$ (23)
We can additionally decompose
$\begin{split}\operatorname*{Var}_{D_{r},\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]\end{split}$ (24)
This reveals the DP-related variance term
$\mathrm{DPVAR}=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\tilde{s}}[g(x;D_{s})]$ (25)
so we have
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\mathrm{RDV}+\mathrm{DPVAR}+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y].$ (26)
Appendix B Non-i.i.d. Synthetic Data
Here, we consider the case of non-i.i.d synthetic datasets, and allow each synthetic dataset to have a different predictor $g_{i}$. We get a similar decomposition as in Theorem 2.2, but the terms of the decomposition are now averages over the possibly different synthetic data distributions, and there is an additional covariance term. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta_{i}|D_{r})$, $i=1,\dotsc,m$, be potentially non-i.i.d. given the real data $D_{r}$. Let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}^{i}|\theta_{i})$, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})$. Then
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\overline{\mathrm{RDV}}+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y],$ (27)
where
$\mathrm{MSE}$ $=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]$ (28)
$\overline{\mathrm{MV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (29)
$\overline{\mathrm{SDV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta_{i}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (30)
$\mathrm{COV}$ $=\frac{1}{m^{2}}\sum_{i\neq j}\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|D_{r}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\right]$ (31)
$\overline{\mathrm{RDV}}$ $=\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right]$ (32)
$\overline{\mathrm{SDB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)]\right]$ (33)
$\overline{\mathrm{MB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].$ (34)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data and $f_{\theta_{i}}$ is the optimal predictor for the synthetic data generating process given parameters $\theta_{i}$. All random quantities are implicitly conditioned on $x$.
Proof.
The classical bias-variance decomposition gives
$\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]=(f(x)-\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})])^{2}+\operatorname*{Var}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]+\operatorname*{Var}_{y}[y].$ (35)
For the bias,
$\begin{split}&f(x)-\operatorname*{\mathbb{E}}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]\\ &=f(x)-\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x,D_{s}^{i})\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-f_{\theta_{i}}(x)+f_{\theta_{i}}(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{\theta_{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-f_{\theta_{i}}(x))\right]+\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f_{\theta_{i}}(x)-g_{i}(x,D_{s}^{i}))\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}(f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)])\right]+\operatorname*{\mathbb{E}}_{D_{r},\theta_{1:m}}\left[\frac{1}{m}\sum_{i=1}^{m}(f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})])\right]\\ &=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}[f_{\theta_{i}}(x)]\right]+\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].\end{split}$ (36)
For the variance,
$\begin{split}\operatorname*{Var}_{D_{r},D_{s}^{1:m}}[\hat{g}(x;D_{s}^{1:m})]&=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{1:m}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]\\ &=\frac{1}{m^{2}}\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]+\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right],\end{split}$ (37)
$\begin{split}\operatorname*{Var}_{D_{s}^{1:m}|D_{r}}\left[\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})\right]&=\sum_{i=1}^{m}\operatorname*{Var}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]+\sum_{i\neq j}\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|D_{r}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\end{split}$ (38)
and
$\operatorname*{Var}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]=\operatorname*{\mathbb{E}}_{\theta_{i}|D_{r}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]+\operatorname*{Var}_{\theta_{i}|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta}[g_{i}(x,D_{s}^{i})].$ (39)
If the synthetic datasets are identically distributed, but not necessarily independent, and the predictors are identical, Theorem B simplifies to
$\mathrm{MSE}=\frac{1}{m}\mathrm{MV}+\frac{1}{m}\mathrm{SDV}+\left(1-\frac{1}{m}\right)\mathrm{COV}+\mathrm{RDV}+(\mathrm{SDB}+\mathrm{MB})^{2}+\operatorname*{Var}_{y}[y],$ (40)
where
$\mathrm{MSE}$ $=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]$ (41)
$\mathrm{MV}$ $=\operatorname*{\mathbb{E}}_{D_{r},\theta}\operatorname*{Var}_{D_{s}|\theta}[g(x;D_{s})]$ (42)
$\mathrm{SDV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\theta|D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]$ (43)
$\mathrm{COV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\left[\operatorname*{Cov}_{D_{s}^{1},D_{s}^{2}|D_{r}}[g(x;D_{s}^{1}),g(x;D_{s}^{2})]\right]$ (44)
$\mathrm{RDV}$ $=\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{D_{s}|D_{r}}[g(x;D_{s})]$ (45)
$\mathrm{SDB}$ $=\operatorname*{\mathbb{E}}_{D_{r}}[f(x)-\operatorname*{\mathbb{E}}_{\theta|D_{r}}[f_{\theta}(x)]]$ (46)
$\mathrm{MB}$ $=\operatorname*{\mathbb{E}}_{D_{r},\theta}[f_{\theta}(x)-\operatorname*{\mathbb{E}}_{D_{s}|\theta}[g(x;D_{s})]].$ (47)
which is Theorem 2.2 with the additional covariance term.
In the noisy summary case, we get an analogue of Theorem 2.4. {theorem}[] Let the parameters for $m$ generators $\theta_{i}\sim p(\theta_{i}|\tilde{s})$, $i=1,\dotsc,m$, be potentially non-i.i.d. given a DP summary $\tilde{s}$ the real data $D_{r}$, let the synthetic datasets be $D_{s}^{i}\sim p(D_{s}^{i}|\theta_{i})$, and let $\hat{g}(x;D_{s}^{1:m})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x;D_{s}^{i})$. Then
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\overline{\mathrm{RDV}}+\overline{\mathrm{DPVAR}}+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y],$ (48)
where
$\mathrm{MSE}$ $=\operatorname*{\mathbb{E}}_{y,D_{r},\tilde{s},D_{s}^{1:m}}[(y-\hat{g}(x;D_{s}^{1:m}))^{2}]$ (49)
$\overline{\mathrm{MV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta_{i}}\operatorname*{Var}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (50)
$\overline{\mathrm{SDV}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\operatorname*{Var}_{\theta_{i}|\tilde{s}}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x;D_{s}^{i})]$ (51)
$\mathrm{COV}$ $=\frac{1}{m^{2}}\sum_{i\neq j}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\left[\operatorname*{Cov}_{D_{s}^{i},D_{s}^{j}|\tilde{s}}[g_{i}(x;D_{s}^{i}),g_{j}(x;D_{s}^{j})]\right]$ (52)
$\overline{\mathrm{RDV}}$ $=\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right]$ (53)
$\overline{\mathrm{DPVAR}}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]$ (54)
$\overline{\mathrm{SDB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s}}\left[f(x)-\operatorname*{\mathbb{E}}_{\theta_{i}|\tilde{s}}[f_{\theta_{i}}(x)]\right]$ (55)
$\overline{\mathrm{MB}}$ $=\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{r},\tilde{s},\theta_{i}}\left[f_{\theta_{i}}(x)-\operatorname*{\mathbb{E}}_{D_{s}^{i}|\theta_{i}}[g_{i}(x,D_{s}^{i})]\right].$ (56)
$f(x)=\operatorname*{\mathbb{E}}_{y}[y]$ is the optimal predictor for real data and $f_{\theta_{i}}$ is the optimal predictor for the synthetic data generating process given parameters $\theta_{i}$. All random quantities are implicitly conditioned on $x$.
Proof.
Using $\tilde{s}$ in place of $D_{r}$ in Theorem B gives
$\mathrm{MSE}=\frac{1}{m}\overline{\mathrm{MV}}+\frac{1}{m}\overline{\mathrm{SDV}}+\mathrm{COV}+\operatorname*{Var}_{D_{r},\tilde{s}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+(\overline{\mathrm{SDB}}+\overline{\mathrm{MB}})^{2}+\operatorname*{Var}_{y}[y].$ (57)
The variance over $D_{r}$ and $\tilde{s}$ can be decomposed
$\begin{split}&\operatorname*{Var}_{D_{r},\tilde{s}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+\operatorname*{Var}_{D_{r}}\operatorname*{\mathbb{E}}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]\\ &=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{Var}_{\tilde{s}|D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|\tilde{s}}[g_{i}(x;D_{s}^{i})]\right]+\operatorname*{Var}_{D_{r}}\left[\frac{1}{m}\sum_{i=1}^{m}\operatorname*{\mathbb{E}}_{D_{s}^{i}|D_{r}}[g_{i}(x;D_{s}^{i})]\right].\end{split}$ (58)
Appendix C Bias-Variance Decomposition for Bregman Divergences
C.1 Background: Bregman Divergences
A Bregman divergence (Bregman, 1967) $D_{F}\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a loss function
$D_{F}(y,g)=F(y)-F(g)-\nabla F(g)^{T}(y-g)$ (59)
where $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is a strictly convex differentiable function. Many common error metrics, like MSE and cross entropy, can be expressed as expected values of a Bregman divergence. In fact, proper scoring rules555 Proper scoring rules are error metrics that are minimised by predicting the correct probabilities. can be characterised via Bregman divergences (Gneiting & Raftery, 2007; Kimpara et al., 2023). Table S1 shows how the metrics we consider are expressed as Bregman divergences (Gupta et al., 2022).
Pfau (2013) derive the following bias-variance decomposition for Bregman divergences:
$\underbrace{\operatorname*{\mathbb{E}}[D(y,g)]}_{\mathrm{Error}}=\underbrace{\operatorname*{\mathbb{E}}[D(y,\operatorname*{\mathbb{E}}y)]}_{\mathrm{Noise}}+\underbrace{D(\operatorname*{\mathbb{E}}y,\mathcal{E}g)}_{\mathrm{Bias}}+\underbrace{\operatorname*{\mathbb{E}}D(\mathcal{E}g,g)}_{\mathrm{Variance}}$ (60)
$y$ is the true value, and $g$ is the predicted value. All of the random quantities are conditioned on $x$. $\mathcal{E}$ is a central prediction:
$\mathcal{E}g=\operatorname*{argmin}_{z}\operatorname*{\mathbb{E}}D(z,g).$ (61)
The variance term can be used to define a generalisation of variance:
$\mathcal{V}g=\operatorname*{\mathbb{E}}D(\mathcal{E}g,g)$ (62)
$\mathcal{E}$ and $\mathcal{V}$ can also be defined conditionally on some random variable $Z$ by making the expectations conditional on $Z$ in the definitions. These obey generalised laws of total expectation and variance (Gupta et al., 2022):
$\mathcal{E}g=\mathcal{E}_{Z}[\mathcal{E}_{g|Z}[g]]$ (63)
and
$\mathcal{V}g=\operatorname*{\mathbb{E}}_{Z}[\mathcal{V}_{g|Z}[g]]+\mathcal{V}_{Z}[\mathcal{E}_{g|Z}[g]].$ (64)
The convex dual of $g$ is $g^{*}=\nabla F(g)$. The central prediction $\mathcal{E}g$ can also be expressed as an expectation over the convex dual (Gupta et al., 2022):
$\mathcal{E}g=(\operatorname*{\mathbb{E}}g^{*})^{*}$ (65)
Table S1: Common error metrics as Bregman divergences. $g$ denotes a prediction in regression and $p$ denotes predicted class probabilities in classification. $g^{(j)}$ and $p^{(j)}$ denote the predictions of different ensemble members. $y$ is the correct value in regression, and a one-hot encoding of the correct class in classification. The binary classification Brier score only looks at probabilities for one class. If the multiclass Brier score is used with two classes, it is twice the binary Brier score.
Error Metric $D_{F}$ $F(t)$ Dual Average
MSE $(y-g)^{2}$ $t^{2}$ $\frac{1}{m}\sum_{j=1}^{m}g^{(j)}$
Brier Score (2 classes) $(y_{1}-p_{1})^{2}$ $t^{2}$ $\frac{1}{m}\sum_{j=1}^{m}p^{(j)}$
Brier Score (Multiclass) $\sum_{i}(y_{i}-p_{i})^{2}$ $\sum_{i}t_{i}^{2}$ $\frac{1}{m}\sum_{j=1}^{m}p^{(j)}$
Cross Entropy $-\sum_{i}y_{i}\ln p_{i}$ $\sum_{i}t_{i}\ln t_{i}$ $\mathrm{softmax}\left(\frac{1}{m}\sum_{j=1}^{m}\ln p^{(j)}\right)$
Gupta et al. (2022) study the bias-variance decomposition of Bregman divergence on a generic ensemble. They show that if the ensemble aggregates prediction by averaging them, bias is not preserved, and can increase. As a solution, they consider dual averaging, that is
$\hat{g}=\left(\frac{1}{m}\sum_{i=1}^{m}g_{i}^{*}\right)^{*}$ (66)
for models $g_{1},\dotsc,g_{m}$ forming the ensemble $\hat{g}$. They show that the bias is preserved in the dual averaged ensemble, and derive a bias-variance decomposition for them. For mean squared error, the dual average is simply the standard average, but for cross entropy, it corresponds to averaging log probabilities.
C.2 Bregman Divergence Decomposition for Synthetic Data
We extend the Bregman divergence decomposition for ensembles from Gupta et al. (2022) to generative ensembles. To prove Theorem C.2, we use the following lemma. {lemma}[Gupta et al. 2022, Proposition 5.3] Let $X_{1},\dotsc,X_{m}$ be i.i.d.  random variables and let $\hat{X}=(\sum_{i=1}^{m}X_{i}^{*})^{*}$ be their dual average. Then $\mathcal{E}\hat{X}=\mathcal{E}X$, $\mathcal{V}\hat{X}\leq\mathcal{V}X$ and for any independent $Y$, $D(\operatorname*{\mathbb{E}}Y,\mathcal{E}\hat{X})=D(\operatorname*{\mathbb{E}}Y,\mathcal{E}X)$.
{theorem}
[] When the synthetic datasets $D_{s}^{1:m}$ are i.i.d.  given the real data $D_{r}$ and $\hat{g}(x;D_{s}^{1:m})=(\frac{1}{m}\sum_{i=1}^{m}g(x;D_{s}^{i})^{*})^{*}$,
$\mathrm{Error}\leq\mathrm{MV}+\mathrm{SDV}+\mathrm{RDV}+\mathrm{Bias}+\mathrm{Noise}$ (67)
where
$\mathrm{Error}$ $=\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[D(y,\hat{g})]$ (68)
$\mathrm{MV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\operatorname*{\mathbb{E}}_{\theta|D_{r}}\mathcal{V}_{D_{s}|\theta}[g]$ (69)
$\mathrm{SDV}$ $=\operatorname*{\mathbb{E}}_{D_{r}}\mathcal{V}_{\theta|D_{r}}\mathcal{E}_{D_{s}|\theta}[g]$ (70)
$\mathrm{RDV}$ $=\mathcal{V}_{D_{r}}\mathcal{E}_{D_{s}|D_{r}}[g]$ (71)
$\mathrm{Bias}$ $=D\left(\operatorname*{\mathbb{E}}_{y}y,\mathcal{E}_{D_{r}}\mathcal{E}_{D_{s}|D_{r}}[g]\right)$ (72)
$\mathrm{Noise}$ $=\operatorname*{\mathbb{E}}_{y}\left[D(y,\operatorname*{\mathbb{E}}_{y}y)\right]$ (73)
Proof.
Plugging the ensemble $\hat{g}$ into the decomposition (60) gives
$\operatorname*{\mathbb{E}}_{y,D_{r},D_{s}^{1:m}}[D(y,\hat{g})]=\operatorname*{\mathbb{E}}_{y}[D(y,\operatorname*{\mathbb{E}}_{y}y)]+D(\operatorname*{\mathbb{E}}_{y}y,\mathcal{E}_{D_{r},D_{s}^{1:m}}\hat{g})+\mathcal{V}_{D_{r},D_{s}^{1:m}}[\hat{g}]$ (74)
Applying the generalised laws of expectation and variance, and Lemma C.2 to the variance term, we obtain:
$\mathcal{V}_{D_{r},D_{s}^{1:m}}[\hat{g}]$ $=\operatorname*{\mathbb{E}}_{D_{r}}\mathcal{V}_{D_{s}^{1:m}|D_{r}}[\hat{g}]+\mathcal{V}_{D_{r}}\mathcal{E}_{D_{s}^{1:m}|D_{r}}[\hat{g}]$ (75)
For the second term on the right:
$\mathcal{E}_{D_{s}^{1:m}|D_{r}}[\hat{g}]=\mathcal{E}_{D_{s}|D_{r}}[g],$ (76)
which gives the RDV:
$\mathcal{V}_{D_{r}}\mathcal{E}_{D_{s}^{1:m}|D_{r}}[\hat{g}]=\mathcal{V}_{D_{r}}\mathcal{E}_{D_{s}|D_{r}}[g].$ (77)
For the first term on the right:
$\operatorname*{\mathbb{E}}_{D_{r}}\mathcal{V}_{D_{s}^{1:m}|D_{r}}[\hat{g}]\leq\operatorname*{\mathbb{E}}_{D_{r}}\mathcal{V}_{D_{s}|D_{r}}[g]$ (78)
and
$\mathcal{V}_{D_{s}|D_{r}}[g]=\operatorname*{\mathbb{E}}_{\theta|D_{r}}\mathcal{V}_{D_{s}|\theta}[g]+\mathcal{V}_{\theta|D_{r}}\mathcal{E}_{D_{s}|\theta}[g],$ (79)
which give MV and SDV.
For the bias
$\begin{split}D(\operatorname*{\mathbb{E}}_{y}y,\mathcal{E}_{D_{r},D_{s}^{1:m}}[\hat{g}])&=D(\operatorname*{\mathbb{E}}_{y}y,\mathcal{E}_{D_{r}}\mathcal{E}_{D_{s}^{1:m}|D_{r}}[\hat{g}])\\ &=D(\operatorname*{\mathbb{E}}_{y}y,\mathcal{E}_{D_{r}}\mathcal{E}_{D_{s}|D_{r}}[g]).\end{split}$ (80)
Putting everything together proves the claim. ∎
This decomposition is not as easy to interpret as the other two in Section 2, as is only gives an upper bound, and does not explicitly depend on the number of synthetic datasets.
Appendix D Experimental Details
D.1 Datasets
In our experiments, we use 7 tabular datasets. For four of them, the downstream prediction task is regression, and for the other three, the prediction task is binary classification. Table S2 lists some general information on the datasets. We use 25% of the real data as a test set, with the remaining 75% being used to generate the synthetic data, for all of the datasets. All experiments are repeated several times, with different train-test splits for each repeat.
Table S2: Details on the datasets used in the experiments. # Cat. and # Num. are the numbers of categorical and numerical features, not counting the target variable. For datasets with removed rows, the table shows the number of rows after the removals.
Dataset # Rows # Cat. # Num. Task
Abalone 4177 1 7 Regression
ACS 2018 50000 5 2 Regression
Adult 45222 8 4 Classification
Reduced Adult 46043 7 2 Classification
Breast Cancer 569 0 30 Classification
California Housing 20622 0 8 Regression
German Credit 1000 13 7 Classification
Insurance 1338 3 3 Regression
Abalone
(Nash et al., 1995, CC BY 4.0) The abalone dataset contains information on abalones, with the task of predicting the number of rings on the abalone from the other information.
ACS 2018
(https://www.census.gov/programs-surveys/acs/microdata/documentation.2018.html, license: https://www.census.gov/data/developers/about/terms-of-service.html) This dataset contains several variables from the American community survey (ACS) of 2018, with the task of predicting a person’s income from the other features. Specifically, the variables we selected are AGEP (age), COW (employer type), SCHL (education), MAR (marital status), WKHP (working hours), SEX, RAC1P (race), and the target PINCP (income). We take a subset of 50000 datapoints from the California data, and log-transform the target variable. We used the folktables package (Ding et al., 2021) to download the given subset of the data.
Adult
(Kohavi & Becker, 1996, CC BY 4.0) The UCI Adult dataset contains general information on people, with the task of predicting whether their income is over \$50000. We drop rows with any missing values.
Reduced Adult
(Kohavi & Becker, 1996, CC BY 4.0) This dataset is the UCI Adult dataset with a reduced set of features. The subset of features is age, workclass, education, marital-status, race, gender, capital-gain, capital-loss and hours-per-week. [^6] This subset was used in the NAPSU-MQ experiments of Räisä et al. (2023b) We binarise capital-gain and capital-loss to indicate whether the original value is positive or not for both synthetic data generation and downstream prediction. We discretise age and hours-per-week to 5 categories for synthetic data generation, and converted back to continuous values for the downstream prediction, as our synthetic data generators require discrete data. We only remove rows with missing values in the included columns, so the number of rows is larger with the reduced set of features.
Breast Cancer
(Wolberg et al., 1995, CC BY 4.0) The breast cancer dataset contains features derived from images of potential tumors, with the task of predicting whether the potential tumor is benign or malignant.
California Housing
(https://scikit-learn.org/stable/datasets/real_world.html#california-housing-dataset, license not mentioned) The california housing dataset contains information on housing districts, specifically census block groups, in California. The task is predicting the median house value in the district. We removed outlier rows where the average number of rooms is at least 50, or the average occupancy is at least 30. According to the dataset description, these likely correspond to districts with many empty houses. We log-transformed the target variable, as well as the population and median income features.
German Credit
(Hofmann, 1994, CC BY 4.0) The German credit dataset contains information on a bank’s customers, with the task of predicting whether the customers are “good” or “bad”.
Insurance
(https://www.kaggle.com/datasets/mirichoi0218/insurance/data, Database Contents License (DbCL) v1.0) The insurance dataset contains general information on people, like age, gender and BMI, as well as the amount they charged their medical insurance, which is the variable to predict. We take a log transform of the target variable before generating synthetic data.
D.2 Downstream Prediction Algorithms
We use the scikit-learn [^7] https://scikit-learn.org/stable/index.html, BSD 3-Clause License implementations of all of the downstream algorithms, which includes probability predictions for all algorithms on the classification tasks. We standardise the data before training for all downstream algorithms except the tree-based algorithms, specifically decision tree, random forest, and gradient boosted trees. This standardisation is done just before downstream training, so the input to the synthetic data generation algorithms is not standardised. We use the default hyperparameters of scikit-learn for all downstream algorithms except MLP, where we increased the maximum number of iterations to 1000, as the default was not enough to converge on some datasets. In particular, this means that decision trees are trained to interpolate the training data, resulting in high variance of the predictions.
D.3 DP Experiment
Both AIM (McKenna et al., 2022) and NAPSU-MQ (Räisä et al., 2023b) generate synthetic data based on noisy values of marginal queries on the real data, which count how many rows of the data have given values for given variables. AIM includes a mechanism that chooses a subset of queries to measure under DP from a potentially very large workload. We set the workload to be all marginal queries over two variables. NAPSU-MQ does not include a query selection mechanism, so we first run AIM with $\epsilon=0.5$, $\delta=\frac{1}{2}n^{-2}$ to select a subset of marginals and then run NAPSU-MQ with the selected marginals with $\epsilon=1$, $\delta=\frac{1}{2}n^{-2}$, which results in the same privacy bounds ($\epsilon=1.5$, $\delta=n^{-2}$; Dwork & Roth, 2014) that we used for AIM. We split the privacy budget between the $m$ synthetic datasets in AIM by dividing the zero-concentrated DP (Bun & Steinke, 2016) parameter that AIM uses internally by $m$, which keeps the total privacy budget fixed.
[…]
D.4 Computational Resources
We ran the synthetic data generation algorithms DDPM, TVAE and CTGAN on a single GPU, synthpop, AIM and NAPSU-MQ on CPU, and all downstream analysis on CPU, all in a cluster environment.
Appendix E Extra Results
E.1 Synthetic Data Generation Algorithms
We compare several synthetic data generation algorithms to see which algorithms are most interesting for subsequent experiments. We use the California housing dataset, where the downstream task is regression. The algorithms we compare are DDPM (Kotelnikov et al., 2023), TVAE (Xu et al., 2019), CTGAN (Xu et al., 2019) and synthpop (Nowok et al., 2016). DDPM, TVAE and CTGAN are a diffusion model, a variational autoencoder and a GAN that are designed for tabular data. We use the implementations from the synthcity library [^11] https://github.com/vanderschaarlab/synthcity, Apache-2.0 license for these. Synthpop generates synthetic data by sampling one column from the real data, and generating the other columns by sequentially training a predictive model on the real data, and predicting the next column from the already generated ones. We use the implementation from the authors (Nowok et al., 2016). [^12] License: GPL-3
We use the default hyperparameters for all of the algorithms. Synthpop and DDPM have a setting that could potentially affect the randomness in the synthetic data generation, so we include both possibilities for these settings in this experiment. For synthpop, this setting is whether the synthetic data is generated from a Bayesian posterior predictive distribution, which synthpop calls “proper” synthetic data. For DDPM, this setting is whether the loss function is MSE or KL divergence. In the plots, the two variants of synthpop are called “SP-P” and “SP-IP” for the proper and improper variants, and the variants of DDPM are “DDPM” and “DDPM-KL”.
The results of the comparison are shown in Figure S1. We see that synthpop and DDPM with MSE loss generally outperform the other generation algorithms, so we select them for the subsequent experiments. There is very little difference between the two variants of synthpop, so we choose the “proper” variant due to its connection with the Bayesian reasoning for using multiple synthetic datasets.
Figure S1: Comparison of synthetic data generation algorithms for several prediction algorithms on the California housing dataset, with 1 to 10 synthetic datasets. DDPM and synthpop achieve smaller MSE in the downstream predictions, so they were selected for further experiments. SP-P and SP-IP are the proper and improper variants of synthpop, and DDPM-KL is DDPM with KL divergence loss. 1-NN and 5-NN are nearest neighbours with 1 and 5 neighbours. The dashed black lines show the performance of each prediction algorithm on the real data, and the solid black line shows the performance of the best predictor, random forest, on the real data. The results are averaged over 3 repeats, with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Linear regression was omitted, as it had nearly identical results as ridge regression. Table S3 in the Appendix contains the numbers in the plots, including ridge regression.
E.2 Extra Plots
Figure S2: MSE on regression datasets of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from DDPM. Increasing the number of synthetic datasets generally decreases MSE, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and ridge regression (RR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. We omitted linear regression from the plots, as it had almost identical results to ridge regression. Tables S4 to S7 contain the numbers from the plots, including linear regression.
Figure S3: All error metrics on the Adult dataset. Note the logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
Figure S4: All error metrics on the breast cancer dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
Figure S5: All error metrics on the German credit dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
(a) Abalone
(b) ACS 2018
Figure S6: MSE prediction on the first two regression datasets. The linear regression measured MSE line for synthpop with ACS 2018 data is cut off due to excluding repeats with extremely large MSE ($\geq 10^{6}$). 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
(a) California Housing
(b) Insurance
Figure S7: MSE prediction on the last two regression datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
(a) Adult
(b) Breast Cancer
Figure S8: Brier score prediction on two classification datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
(a) Regression
(b) Classification
Figure S9: Random forest performance prediction on the regression datasets in (a) and classification datasets in (b). The prediction is reasonably accurate on the datasets with accurate estimates of the error. On the other datasets, the prediction can have high variance. The lines show averages over 3 different train-test splits and 3 repeats of model training per split. The error bands are 95% confidence intervals formed by bootsrapping over the repeats and different splits.
Figure S10: All error metrics on the Adult dataset with reduced features and DP synthetic data generation. Note the he logarithmic scale on the cross entropy y-axis. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
E.3 Result Tables
Table S3: Table of synthetic data generator comparison results from Figure S1. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
m 1 2 5 10
Downstream Generator
1-NN CTGAN 0.2621 $\pm$ 0.0228 0.1775 $\pm$ 0.0037 0.1316 $\pm$ 0.0044 0.1164 $\pm$ 0.0040
DDPM 0.1365 $\pm$ 0.0029 0.1365 $\pm$ 0.0029 0.1365 $\pm$ 0.0029 0.1343 $\pm$ 0.0019
DDPM-KL 0.2447 $\pm$ 0.0082 0.2447 $\pm$ 0.0082 0.2238 $\pm$ 0.0240 0.2124 $\pm$ 0.0085
SP-IP 0.1544 $\pm$ 0.0040 0.1139 $\pm$ 0.0028 0.0892 $\pm$ 0.0018 0.0805 $\pm$ 0.0019
SP-P 0.1654 $\pm$ 0.0014 0.1172 $\pm$ 0.0012 0.0893 $\pm$ 0.0027 0.0801 $\pm$ 0.0015
TVAE 0.2195 $\pm$ 0.0036 0.2002 $\pm$ 0.0316 0.2114 $\pm$ 0.0123 0.1829 $\pm$ 0.0073
5-NN CTGAN 0.1518 $\pm$ 0.0110 0.1230 $\pm$ 0.0047 0.1125 $\pm$ 0.0050 0.1092 $\pm$ 0.0040
DDPM 0.0885 $\pm$ 0.0023 0.0885 $\pm$ 0.0023 0.0885 $\pm$ 0.0023 0.0881 $\pm$ 0.0020
DDPM-KL 0.1379 $\pm$ 0.0015 0.1379 $\pm$ 0.0015 0.1342 $\pm$ 0.0023 0.1320 $\pm$ 0.0023
SP-IP 0.0936 $\pm$ 0.0030 0.0852 $\pm$ 0.0027 0.0797 $\pm$ 0.0021 0.0780 $\pm$ 0.0021
SP-P 0.1029 $\pm$ 0.0013 0.0891 $\pm$ 0.0020 0.0814 $\pm$ 0.0015 0.0788 $\pm$ 0.0014
TVAE 0.1240 $\pm$ 0.0008 0.1201 $\pm$ 0.0073 0.1227 $\pm$ 0.0028 0.1161 $\pm$ 0.0018
Decision Tree CTGAN 0.2926 $\pm$ 0.0398 0.1818 $\pm$ 0.0048 0.1301 $\pm$ 0.0034 0.1136 $\pm$ 0.0024
DDPM 0.1332 $\pm$ 0.0076 0.1285 $\pm$ 0.0068 0.1266 $\pm$ 0.0069 0.1219 $\pm$ 0.0033
DDPM-KL 0.2395 $\pm$ 0.0041 0.2314 $\pm$ 0.0043 0.2024 $\pm$ 0.0272 0.1885 $\pm$ 0.0186
SP-IP 0.1355 $\pm$ 0.0048 0.0981 $\pm$ 0.0010 0.0762 $\pm$ 0.0014 0.0689 $\pm$ 0.0008
SP-P 0.1432 $\pm$ 0.0058 0.1015 $\pm$ 0.0011 0.0758 $\pm$ 0.0013 0.0669 $\pm$ 0.0010
TVAE 0.2086 $\pm$ 0.0063 0.1838 $\pm$ 0.0328 0.1894 $\pm$ 0.0149 0.1620 $\pm$ 0.0091
Random Forest CTGAN 0.1286 $\pm$ 0.0151 0.1050 $\pm$ 0.0032 0.1004 $\pm$ 0.0026 0.0988 $\pm$ 0.0024
DDPM 0.0681 $\pm$ 0.0020 0.0678 $\pm$ 0.0022 0.0674 $\pm$ 0.0022 0.0673 $\pm$ 0.0021
DDPM-KL 0.1113 $\pm$ 0.0018 0.1106 $\pm$ 0.0019 0.1089 $\pm$ 0.0008 0.1082 $\pm$ 0.0022
SP-IP 0.0676 $\pm$ 0.0012 0.0649 $\pm$ 0.0010 0.0626 $\pm$ 0.0008 0.0619 $\pm$ 0.0007
SP-P 0.0762 $\pm$ 0.0020 0.0696 $\pm$ 0.0014 0.0642 $\pm$ 0.0020 0.0629 $\pm$ 0.0018
TVAE 0.0912 $\pm$ 0.0013 0.0888 $\pm$ 0.0024 0.0889 $\pm$ 0.0017 0.0866 $\pm$ 0.0022
MLP CTGAN 0.1267 $\pm$ 0.0082 0.1054 $\pm$ 0.0105 0.0991 $\pm$ 0.0075 0.0958 $\pm$ 0.0049
DDPM 0.0650 $\pm$ 0.0014 0.0642 $\pm$ 0.0016 0.0638 $\pm$ 0.0007 0.0635 $\pm$ 0.0009
DDPM-KL 0.1032 $\pm$ 0.0020 0.1020 $\pm$ 0.0028 0.0999 $\pm$ 0.0023 0.0985 $\pm$ 0.0036
SP-IP 0.0702 $\pm$ 0.0018 0.0682 $\pm$ 0.0014 0.0661 $\pm$ 0.0011 0.0658 $\pm$ 0.0007
SP-P 0.0747 $\pm$ 0.0029 0.0705 $\pm$ 0.0006 0.0685 $\pm$ 0.0003 0.0681 $\pm$ 0.0006
TVAE 0.0882 $\pm$ 0.0065 0.0834 $\pm$ 0.0023 0.0818 $\pm$ 0.0019 0.0809 $\pm$ 0.0021
Gradient Boosting CTGAN 0.1270 $\pm$ 0.0162 0.1053 $\pm$ 0.0044 0.1026 $\pm$ 0.0030 0.1012 $\pm$ 0.0023
DDPM 0.0725 $\pm$ 0.0023 0.0725 $\pm$ 0.0023 0.0725 $\pm$ 0.0023 0.0724 $\pm$ 0.0022
DDPM-KL 0.1054 $\pm$ 0.0038 0.1054 $\pm$ 0.0038 0.1045 $\pm$ 0.0031 0.1044 $\pm$ 0.0037
SP-IP 0.0715 $\pm$ 0.0018 0.0698 $\pm$ 0.0013 0.0689 $\pm$ 0.0012 0.0685 $\pm$ 0.0011
SP-P 0.0745 $\pm$ 0.0023 0.0724 $\pm$ 0.0015 0.0711 $\pm$ 0.0012 0.0708 $\pm$ 0.0018
TVAE 0.0912 $\pm$ 0.0026 $0$$\pm$ 0.0018 0.0901 $\pm$ 0.0018 0.0893 $\pm$ 0.0018
SVM CTGAN 0.1235 $\pm$ 0.0144 0.1051 $\pm$ 0.0017 0.1005 $\pm$ 0.0024 0.0993 $\pm$ 0.0022
DDPM $0$$\pm$ 0.0027 $0$$\pm$ 0.0027 $0$$\pm$ 0.0027 $0$$\pm$ 0.0027
DDPM-KL 0.1026 $\pm$ 0.0032 0.1026 $\pm$ 0.0032 0.1017 $\pm$ 0.0024 0.1013 $\pm$ 0.0034
SP-IP 0.0736 $\pm$ 0.0028 0.0725 $\pm$ 0.0026 0.0712 $\pm$ 0.0022 0.0711 $\pm$ 0.0020
SP-P 0.0768 $\pm$ 0.0025 0.0745 $\pm$ 0.0020 0.0728 $\pm$ 0.0018 0.0724 $\pm$ 0.0020
TVAE 0.0944 $\pm$ 0.0022 0.0934 $\pm$ 0.0004 0.0938 $\pm$ 0.0011 0.0919 $\pm$ 0.0023
Ridge Regression CTGAN 0.1587 $\pm$ 0.0182 0.1391 $\pm$ 0.0032 0.1381 $\pm$ 0.0042 0.1372 $\pm$ 0.0044
DDPM 0.1110 $\pm$ 0.0036 0.1110 $\pm$ 0.0036 0.1110 $\pm$ 0.0036 0.1109 $\pm$ 0.0035
DDPM-KL 0.1296 $\pm$ 0.0035 0.1296 $\pm$ 0.0035 0.1295 $\pm$ 0.0035 0.1293 $\pm$ 0.0037
SP-IP 0.1098 $\pm$ 0.0032 0.1095 $\pm$ 0.0031 0.1095 $\pm$ 0.0030 0.1096 $\pm$ 0.0030
SP-P 0.1109 $\pm$ 0.0034 0.1106 $\pm$ 0.0030 0.1105 $\pm$ 0.0030 0.1105 $\pm$ 0.0033
TVAE 0.1353 $\pm$ 0.0044 0.1343 $\pm$ 0.0053 0.1341 $\pm$ 0.0055 0.1336 $\pm$ 0.0053
Linear Regression CTGAN 0.1587 $\pm$ 0.0182 0.1391 $\pm$ 0.0032 0.1381 $\pm$ 0.0042 0.1372 $\pm$ 0.0044
DDPM 0.1110 $\pm$ 0.0036 0.1110 $\pm$ 0.0036 0.1110 $\pm$ 0.0036 0.1109 $\pm$ 0.0035
DDPM-KL 0.1296 $\pm$ 0.0035 0.1296 $\pm$ 0.0035 0.1295 $\pm$ 0.0035 0.1293 $\pm$ 0.0037
SP-IP 0.1098 $\pm$ 0.0032 0.1095 $\pm$ 0.0031 0.1095 $\pm$ 0.0030 0.1096 $\pm$ 0.0030
SP-P 0.1109 $\pm$ 0.0034 0.1105 $\pm$ 0.0030 0.1105 $\pm$ 0.0030 0.1105 $\pm$ 0.0033
TVAE 0.1352 $\pm$ 0.0044 0.1343 $\pm$ 0.0053 0.1341 $\pm$ 0.0055 0.1336 $\pm$ 0.0053
Table S4: Table of results from the Abalone dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
Linear Regression DDPM 8.24 $\pm$ 2.392 8.24 $\pm$ 2.392 8.31 $\pm$ 2.334 8.25 $\pm$ 2.388 7.79 $\pm$ 1.602 7.79 $\pm$ 1.602
SP-P 5.57 $\pm$ 0.646 5.42 $\pm$ 0.429 5.35 $\pm$ 0.443 5.38 $\pm$ 0.470 5.35 $\pm$ 0.463 5.33 $\pm$ 0.452
Ridge Regression DDPM 8.24 $\pm$ 2.392 8.24 $\pm$ 2.392 8.31 $\pm$ 2.335 8.25 $\pm$ 2.388 7.79 $\pm$ 1.599 7.79 $\pm$ 1.599
SP-P 5.57 $\pm$ 0.648 5.43 $\pm$ 0.432 5.35 $\pm$ 0.445 5.39 $\pm$ 0.472 5.35 $\pm$ 0.465 5.33 $\pm$ 0.454
1-NN DDPM 9.01 $\pm$ 0.621 9.01 $\pm$ 0.621 8.57 $\pm$ 0.621 8.27 $\pm$ 0.123 7.72 $\pm$ 0.434 7.72 $\pm$ 0.434
SP-P 9.52 $\pm$ 0.890 7.35 $\pm$ 0.764 6.21 $\pm$ 0.482 5.66 $\pm$ 0.514 5.30 $\pm$ 0.439 5.13 $\pm$ 0.479
5-NN DDPM 5.76 $\pm$ 0.457 5.76 $\pm$ 0.457 5.71 $\pm$ 0.399 5.66 $\pm$ 0.493 5.56 $\pm$ 0.491 5.56 $\pm$ 0.491
SP-P 6.32 $\pm$ 0.932 5.75 $\pm$ 0.704 5.42 $\pm$ 0.550 5.26 $\pm$ 0.539 5.18 $\pm$ 0.523 5.14 $\pm$ 0.513
Decision Tree DDPM 9.17 $\pm$ 0.199 8.79 $\pm$ 0.289 8.19 $\pm$ 0.741 7.81 $\pm$ 0.509 7.17 $\pm$ 0.174 7.17 $\pm$ 0.194
SP-P 9.37 $\pm$ 0.523 7.38 $\pm$ 0.418 6.05 $\pm$ 0.237 5.43 $\pm$ 0.165 5.07 $\pm$ 0.274 4.95 $\pm$ 0.348
Random Forest DDPM 5.36 $\pm$ 0.372 5.34 $\pm$ 0.405 5.24 $\pm$ 0.361 5.21 $\pm$ 0.399 5.14 $\pm$ 0.360 5.13 $\pm$ 0.361
SP-P 5.76 $\pm$ 0.378 5.40 $\pm$ 0.282 5.12 $\pm$ 0.268 4.96 $\pm$ 0.253 4.87 $\pm$ 0.276 4.85 $\pm$ 0.318
Gradient Boosting DDPM 5.47 $\pm$ 0.506 5.47 $\pm$ 0.506 5.47 $\pm$ 0.503 5.45 $\pm$ 0.526 5.39 $\pm$ 0.509 5.39 $\pm$ 0.508
SP-P 5.45 $\pm$ 0.336 5.25 $\pm$ 0.301 5.08 $\pm$ 0.294 4.99 $\pm$ 0.329 4.96 $\pm$ 0.324 4.96 $\pm$ 0.363
MLP DDPM 5.06 $\pm$ 0.323 5.04 $\pm$ 0.335 5.00 $\pm$ 0.249 5.00 $\pm$ 0.273 4.92 $\pm$ 0.267 4.92 $\pm$ 0.280
SP-P 6.02 $\pm$ 1.935 5.38 $\pm$ 1.075 4.96 $\pm$ 0.550 4.97 $\pm$ 0.570 4.84 $\pm$ 0.399 4.81 $\pm$ 0.399
SVM DDPM 5.55 $\pm$ 0.592 5.55 $\pm$ 0.592 5.55 $\pm$ 0.586 5.54 $\pm$ 0.598 5.51 $\pm$ 0.610 5.51 $\pm$ 0.610
SP-P 5.38 $\pm$ 0.515 5.35 $\pm$ 0.546 5.28 $\pm$ 0.499 5.28 $\pm$ 0.583 5.27 $\pm$ 0.604 5.26 $\pm$ 0.588
Table S5: Table of results from the ACS 2018 dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats. The nans for linear regression are caused by excluding repeats with extremely large MSE ($\geq 10^{6}$).
m 1 2 4 8 16 32
Downstream Generator
Linear Regression DDPM 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.005 0.87 $\pm$ 0.001 0.87 $\pm$ 0.003 0.87 $\pm$ 0.003
SP-P 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.007 0.78 $\pm$ nan nan nan
Ridge Regression DDPM 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.005 0.87 $\pm$ 0.003 0.87 $\pm$ 0.004 0.87 $\pm$ 0.003
SP-P 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.007 0.78 $\pm$ 0.007 0.77 $\pm$ 0.007 0.77 $\pm$ 0.006
1-NN DDPM 1.55 $\pm$ 0.022 1.26 $\pm$ 0.026 1.10 $\pm$ 0.019 1.03 $\pm$ 0.014 1.00 $\pm$ 0.013 0.98 $\pm$ 0.011
SP-P 1.38 $\pm$ 0.015 1.04 $\pm$ 0.010 0.88 $\pm$ 0.001 0.79 $\pm$ 0.004 0.75 $\pm$ 0.006 0.73 $\pm$ 0.007
5-NN DDPM 0.98 $\pm$ 0.006 0.93 $\pm$ 0.009 0.90 $\pm$ 0.008 0.89 $\pm$ 0.007 0.89 $\pm$ 0.004 0.88 $\pm$ 0.004
SP-P 0.93 $\pm$ 0.002 0.82 $\pm$ 0.005 0.78 $\pm$ 0.004 0.75 $\pm$ 0.008 0.74 $\pm$ 0.008 0.73 $\pm$ 0.009
Decision Tree DDPM 1.41 $\pm$ 0.008 1.09 $\pm$ 0.005 0.95 $\pm$ 0.006 0.87 $\pm$ 0.004 0.84 $\pm$ 0.006 0.82 $\pm$ 0.011
SP-P 1.25 $\pm$ 0.035 0.96 $\pm$ 0.023 0.82 $\pm$ 0.011 0.74 $\pm$ 0.005 0.71 $\pm$ 0.003 0.69 $\pm$ 0.004
Random Forest DDPM 0.88 $\pm$ 0.010 0.82 $\pm$ 0.013 0.80 $\pm$ 0.009 0.78 $\pm$ 0.009 0.78 $\pm$ 0.009 0.77 $\pm$ 0.009
SP-P 0.84 $\pm$ 0.010 0.76 $\pm$ 0.010 0.72 $\pm$ 0.006 0.69 $\pm$ 0.005 0.68 $\pm$ 0.005 0.68 $\pm$ 0.006
Gradient Boosting DDPM 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.006 0.74 $\pm$ 0.005 0.74 $\pm$ 0.004 0.74 $\pm$ 0.004
SP-P 0.68 $\pm$ 0.005 0.68 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.006 0.67 $\pm$ 0.006
MLP DDPM 0.82 $\pm$ 0.019 0.77 $\pm$ 0.018 0.76 $\pm$ 0.017 0.75 $\pm$ 0.007 0.74 $\pm$ 0.005 0.74 $\pm$ 0.004
SP-P 0.76 $\pm$ 0.006 0.71 $\pm$ 0.003 0.68 $\pm$ 0.008 0.67 $\pm$ 0.007 0.67 $\pm$ 0.007 0.66 $\pm$ 0.006
SVM DDPM 0.79 $\pm$ 0.004 0.78 $\pm$ 0.002 0.77 $\pm$ 0.002 0.78 $\pm$ 0.002 0.78 $\pm$ 0.003 0.78 $\pm$ 0.002
SP-P 0.71 $\pm$ 0.005 0.69 $\pm$ 0.003 0.69 $\pm$ 0.002 0.69 $\pm$ 0.004 0.69 $\pm$ 0.004 0.68 $\pm$ 0.004
Table S6: Table of results from the California Housing dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
Linear Regression DDPM 0.12 $\pm$ 0.004 0.12 $\pm$ 0.004 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003
SP-P 0.12 $\pm$ 0.001 0.11 $\pm$ 0.002 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.002 0.11 $\pm$ 0.002
Ridge Regression DDPM 0.12 $\pm$ 0.004 0.12 $\pm$ 0.004 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003 0.12 $\pm$ 0.003
SP-P 0.12 $\pm$ 0.001 0.11 $\pm$ 0.002 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.002 0.11 $\pm$ 0.002
1-NN DDPM 0.14 $\pm$ 0.008 0.14 $\pm$ 0.008 0.12 $\pm$ 0.006 0.13 $\pm$ 0.007 0.12 $\pm$ 0.004 0.12 $\pm$ 0.004
SP-P 0.16 $\pm$ 0.002 0.12 $\pm$ 0.002 0.09 $\pm$ 0.002 0.08 $\pm$ 0.001 0.08 $\pm$ 0.002 0.08 $\pm$ 0.001
5-NN DDPM 0.09 $\pm$ 0.002 0.09 $\pm$ 0.002 0.09 $\pm$ 0.002 0.09 $\pm$ 0.002 0.09 $\pm$ 0.002 0.09 $\pm$ 0.002
SP-P 0.10 $\pm$ 0.003 0.09 $\pm$ 0.001 0.08 $\pm$ 0.001 0.08 $\pm$ 0.001 0.08 $\pm$ 0.001 0.08 $\pm$ 0.001
Decision Tree DDPM 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.10 $\pm$ 0.004 0.11 $\pm$ 0.006 0.10 $\pm$ 0.001 0.10 $\pm$ 0.002
SP-P 0.15 $\pm$ 0.004 0.10 $\pm$ 0.004 0.08 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.06 $\pm$ 0.001
Random Forest DDPM 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001
SP-P 0.08 $\pm$ 0.002 0.07 $\pm$ 0.003 0.07 $\pm$ 0.002 0.06 $\pm$ 0.001 0.06 $\pm$ 0.001 0.06 $\pm$ 0.001
Gradient Boosting DDPM 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002
SP-P 0.08 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001
MLP DDPM 0.07 $\pm$ 0.000 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.06 $\pm$ 0.001 0.06 $\pm$ 0.001
SP-P 0.08 $\pm$ 0.001 0.07 $\pm$ 0.002 0.07 $\pm$ 0.002 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001
SVM DDPM 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001
SP-P 0.08 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001 0.07 $\pm$ 0.001
Table S7: Table of results from the Insurance dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
Linear Regression DDPM 0.25 $\pm$ 0.037 0.24 $\pm$ 0.039 0.23 $\pm$ 0.030 0.23 $\pm$ 0.023 0.23 $\pm$ 0.020 0.23 $\pm$ 0.021
SP-P 0.20 $\pm$ 0.009 0.20 $\pm$ 0.009 0.20 $\pm$ 0.008 0.20 $\pm$ 0.006 0.20 $\pm$ 0.006 0.20 $\pm$ 0.006
Ridge Regression DDPM 0.25 $\pm$ 0.037 0.24 $\pm$ 0.039 0.23 $\pm$ 0.031 0.23 $\pm$ 0.023 0.23 $\pm$ 0.021 0.23 $\pm$ 0.021
SP-P 0.20 $\pm$ 0.009 0.20 $\pm$ 0.010 0.20 $\pm$ 0.008 0.20 $\pm$ 0.007 0.20 $\pm$ 0.006 0.20 $\pm$ 0.006
1-NN DDPM 0.40 $\pm$ 0.076 0.37 $\pm$ 0.104 0.34 $\pm$ 0.056 0.33 $\pm$ 0.042 0.34 $\pm$ 0.040 0.34 $\pm$ 0.040
SP-P 0.29 $\pm$ 0.063 0.20 $\pm$ 0.012 0.17 $\pm$ 0.015 0.15 $\pm$ 0.009 0.15 $\pm$ 0.006 0.14 $\pm$ 0.002
5-NN DDPM 0.23 $\pm$ 0.031 0.21 $\pm$ 0.031 0.21 $\pm$ 0.017 0.20 $\pm$ 0.009 0.20 $\pm$ 0.014 0.21 $\pm$ 0.017
SP-P 0.19 $\pm$ 0.014 0.16 $\pm$ 0.007 0.15 $\pm$ 0.008 0.15 $\pm$ 0.005 0.15 $\pm$ 0.005 0.14 $\pm$ 0.003
Decision Tree DDPM 0.38 $\pm$ 0.074 0.34 $\pm$ 0.110 0.32 $\pm$ 0.075 0.31 $\pm$ 0.071 0.32 $\pm$ 0.070 0.33 $\pm$ 0.069
SP-P 0.26 $\pm$ 0.011 0.21 $\pm$ 0.012 0.17 $\pm$ 0.003 0.16 $\pm$ 0.002 0.15 $\pm$ 0.007 0.14 $\pm$ 0.007
Random Forest DDPM 0.22 $\pm$ 0.015 0.20 $\pm$ 0.028 0.19 $\pm$ 0.015 0.19 $\pm$ 0.013 0.19 $\pm$ 0.009 0.19 $\pm$ 0.010
SP-P 0.18 $\pm$ 0.023 0.16 $\pm$ 0.019 0.15 $\pm$ 0.007 0.14 $\pm$ 0.006 0.14 $\pm$ 0.008 0.14 $\pm$ 0.007
Gradient Boosting DDPM 0.18 $\pm$ 0.017 0.17 $\pm$ 0.011 0.17 $\pm$ 0.004 0.17 $\pm$ 0.005 0.17 $\pm$ 0.013 0.17 $\pm$ 0.016
SP-P 0.16 $\pm$ 0.009 0.15 $\pm$ 0.009 0.14 $\pm$ 0.007 0.14 $\pm$ 0.009 0.14 $\pm$ 0.010 0.13 $\pm$ 0.009
MLP DDPM 0.23 $\pm$ 0.010 0.23 $\pm$ 0.008 0.21 $\pm$ 0.008 0.20 $\pm$ 0.009 0.20 $\pm$ 0.013 0.20 $\pm$ 0.014
SP-P 0.16 $\pm$ 0.009 0.16 $\pm$ 0.011 0.15 $\pm$ 0.009 0.15 $\pm$ 0.009 0.15 $\pm$ 0.011 0.15 $\pm$ 0.012
SVM DDPM 0.18 $\pm$ 0.018 0.17 $\pm$ 0.023 0.16 $\pm$ 0.015 0.16 $\pm$ 0.010 0.16 $\pm$ 0.004 0.16 $\pm$ 0.004
SP-P 0.15 $\pm$ 0.015 0.14 $\pm$ 0.014 0.14 $\pm$ 0.015 0.14 $\pm$ 0.015 0.14 $\pm$ 0.015 0.14 $\pm$ 0.014
Table S8: Table of results from the Adult dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
1-NN DDPM - Log Prob. Avg. 0.24 $\pm$ 0.005 0.19 $\pm$ 0.004 0.19 $\pm$ 0.006 0.18 $\pm$ 0.002 0.18 $\pm$ 0.003 0.17 $\pm$ 0.003
DDPM - Prob. Avg. 0.24 $\pm$ 0.005 0.19 $\pm$ 0.004 0.16 $\pm$ 0.005 0.15 $\pm$ 0.002 0.14 $\pm$ 0.002 0.14 $\pm$ 0.002
SP-P - Log Prob. Avg. 0.21 $\pm$ 0.005 0.16 $\pm$ 0.003 0.15 $\pm$ 0.003 0.15 $\pm$ 0.005 0.14 $\pm$ 0.003 0.14 $\pm$ 0.003
SP-P - Prob. Avg. 0.21 $\pm$ 0.005 0.16 $\pm$ 0.003 0.13 $\pm$ 0.002 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
5-NN DDPM - Log Prob. Avg. 0.14 $\pm$ 0.003 0.14 $\pm$ 0.001 0.14 $\pm$ 0.002 0.14 $\pm$ 0.002 0.13 $\pm$ 0.001 0.13 $\pm$ 0.000
DDPM - Prob. Avg. 0.14 $\pm$ 0.003 0.13 $\pm$ 0.001 0.13 $\pm$ 0.001 0.13 $\pm$ 0.001 0.13 $\pm$ 0.000 0.12 $\pm$ 0.000
SP-P - Log Prob. Avg. 0.14 $\pm$ 0.003 0.14 $\pm$ 0.004 0.14 $\pm$ 0.003 0.13 $\pm$ 0.003 0.13 $\pm$ 0.001 0.13 $\pm$ 0.002
SP-P - Prob. Avg. 0.14 $\pm$ 0.003 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
Decision Tree DDPM - Log Prob. Avg. 0.21 $\pm$ 0.000 0.17 $\pm$ 0.001 0.16 $\pm$ 0.001 0.15 $\pm$ 0.002 0.15 $\pm$ 0.000 0.14 $\pm$ 0.001
DDPM - Prob. Avg. 0.21 $\pm$ 0.000 0.16 $\pm$ 0.001 0.14 $\pm$ 0.001 0.13 $\pm$ 0.001 0.12 $\pm$ 0.000 0.12 $\pm$ 0.000
SP-P - Log Prob. Avg. 0.18 $\pm$ 0.003 0.15 $\pm$ 0.001 0.14 $\pm$ 0.002 0.13 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
SP-P - Prob. Avg. 0.18 $\pm$ 0.003 0.14 $\pm$ 0.001 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
Random Forest DDPM - Log Prob. Avg. 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
DDPM - Prob. Avg. 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
SP-P - Log Prob. Avg. 0.12 $\pm$ 0.002 0.11 $\pm$ 0.002 0.11 $\pm$ 0.002 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
SP-P - Prob. Avg. 0.12 $\pm$ 0.002 0.11 $\pm$ 0.002 0.10 $\pm$ 0.002 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
MLP DDPM - Log Prob. Avg. 0.14 $\pm$ 0.006 0.13 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
DDPM - Prob. Avg. 0.14 $\pm$ 0.006 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
SP-P - Log Prob. Avg. 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001
SP-P - Prob. Avg. 0.13 $\pm$ 0.003 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
Gradient Boosting DDPM - Log Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
DDPM - Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
SP-P - Log Prob. Avg. 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
SP-P - Prob. Avg. 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
SVM DDPM - Log Prob. Avg. 0.12 $\pm$ 0.003 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.000 0.12 $\pm$ 0.000 0.12 $\pm$ 0.001
DDPM - Prob. Avg. 0.12 $\pm$ 0.003 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.000 0.12 $\pm$ 0.000 0.12 $\pm$ 0.001
SP-P - Log Prob. Avg. 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
SP-P - Prob. Avg. 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
Logistic Regression DDPM - Log Prob. Avg. 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000
DDPM - Prob. Avg. 0.12 $\pm$ 0.001 0.11 $\pm$ 0.002 0.12 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.000 0.11 $\pm$ 0.000
SP-P - Log Prob. Avg. 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
SP-P - Prob. Avg. 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001 0.10 $\pm$ 0.001
Table S9: Table of results from the Breast Cancer dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
1-NN DDPM - Log Prob. Avg. 0.12 $\pm$ 0.075 0.09 $\pm$ 0.054 0.09 $\pm$ 0.054 0.10 $\pm$ 0.043 0.10 $\pm$ 0.043 0.10 $\pm$ 0.043
DDPM - Prob. Avg. 0.12 $\pm$ 0.075 0.09 $\pm$ 0.054 0.09 $\pm$ 0.054 0.09 $\pm$ 0.049 0.09 $\pm$ 0.048 0.09 $\pm$ 0.047
SP-P - Log Prob. Avg. 0.11 $\pm$ 0.045 0.07 $\pm$ 0.015 0.05 $\pm$ 0.013 0.04 $\pm$ 0.002 0.04 $\pm$ 0.003 0.03 $\pm$ 0.003
SP-P - Prob. Avg. 0.11 $\pm$ 0.045 0.07 $\pm$ 0.015 0.05 $\pm$ 0.011 0.04 $\pm$ 0.004 0.04 $\pm$ 0.000 0.03 $\pm$ 0.001
5-NN DDPM - Log Prob. Avg. 0.14 $\pm$ 0.086 0.13 $\pm$ 0.087 0.13 $\pm$ 0.087 0.13 $\pm$ 0.081 0.12 $\pm$ 0.070 0.12 $\pm$ 0.077
DDPM - Prob. Avg. 0.14 $\pm$ 0.086 0.12 $\pm$ 0.068 0.12 $\pm$ 0.068 0.11 $\pm$ 0.061 0.11 $\pm$ 0.058 0.11 $\pm$ 0.060
SP-P - Log Prob. Avg. 0.05 $\pm$ 0.012 0.04 $\pm$ 0.005 0.03 $\pm$ 0.007 0.03 $\pm$ 0.003 0.03 $\pm$ 0.005 0.02 $\pm$ 0.004
SP-P - Prob. Avg. 0.05 $\pm$ 0.012 0.04 $\pm$ 0.003 0.04 $\pm$ 0.002 0.03 $\pm$ 0.002 0.03 $\pm$ 0.001 0.03 $\pm$ 0.000
Decision Tree DDPM - Log Prob. Avg. 0.14 $\pm$ 0.083 0.09 $\pm$ 0.026 0.09 $\pm$ 0.026 0.10 $\pm$ 0.033 0.10 $\pm$ 0.033 0.10 $\pm$ 0.031
DDPM - Prob. Avg. 0.14 $\pm$ 0.083 0.09 $\pm$ 0.026 0.09 $\pm$ 0.025 0.08 $\pm$ 0.020 0.09 $\pm$ 0.025 0.09 $\pm$ 0.023
SP-P - Log Prob. Avg. 0.10 $\pm$ 0.028 0.05 $\pm$ 0.012 0.04 $\pm$ 0.009 0.03 $\pm$ 0.004 0.03 $\pm$ 0.009 0.03 $\pm$ 0.006
SP-P - Prob. Avg. 0.10 $\pm$ 0.028 0.05 $\pm$ 0.012 0.04 $\pm$ 0.006 0.03 $\pm$ 0.003 0.03 $\pm$ 0.006 0.03 $\pm$ 0.004
Random Forest DDPM - Log Prob. Avg. 0.08 $\pm$ 0.033 0.07 $\pm$ 0.018 0.07 $\pm$ 0.019 0.07 $\pm$ 0.019 0.07 $\pm$ 0.022 0.07 $\pm$ 0.022
DDPM - Prob. Avg. 0.08 $\pm$ 0.033 0.07 $\pm$ 0.018 0.07 $\pm$ 0.019 0.07 $\pm$ 0.019 0.07 $\pm$ 0.022 0.07 $\pm$ 0.022
SP-P - Log Prob. Avg. 0.04 $\pm$ 0.004 0.03 $\pm$ 0.008 0.03 $\pm$ 0.008 0.03 $\pm$ 0.007 0.03 $\pm$ 0.006 0.03 $\pm$ 0.006
SP-P - Prob. Avg. 0.04 $\pm$ 0.004 0.03 $\pm$ 0.007 0.03 $\pm$ 0.007 0.03 $\pm$ 0.006 0.03 $\pm$ 0.006 0.03 $\pm$ 0.006
MLP DDPM - Log Prob. Avg. 0.05 $\pm$ 0.015 0.05 $\pm$ 0.008 0.04 $\pm$ 0.007 0.05 $\pm$ 0.010 0.05 $\pm$ 0.016 0.05 $\pm$ 0.017
DDPM - Prob. Avg. 0.05 $\pm$ 0.015 0.05 $\pm$ 0.007 0.05 $\pm$ 0.005 0.05 $\pm$ 0.009 0.05 $\pm$ 0.015 0.05 $\pm$ 0.016
SP-P - Log Prob. Avg. 0.05 $\pm$ 0.012 0.04 $\pm$ 0.010 0.03 $\pm$ 0.006 0.03 $\pm$ 0.001 0.02 $\pm$ 0.007 0.02 $\pm$ 0.006
SP-P - Prob. Avg. 0.05 $\pm$ 0.012 0.04 $\pm$ 0.009 0.03 $\pm$ 0.003 0.03 $\pm$ 0.002 0.03 $\pm$ 0.004 0.02 $\pm$ 0.003
Gradient Boosting DDPM - Log Prob. Avg. 0.13 $\pm$ 0.089 0.08 $\pm$ 0.031 0.09 $\pm$ 0.034 0.08 $\pm$ 0.029 0.08 $\pm$ 0.033 0.08 $\pm$ 0.032
DDPM - Prob. Avg. 0.13 $\pm$ 0.089 0.09 $\pm$ 0.031 0.09 $\pm$ 0.033 0.08 $\pm$ 0.029 0.08 $\pm$ 0.031 0.08 $\pm$ 0.030
SP-P - Log Prob. Avg. 0.04 $\pm$ 0.002 0.04 $\pm$ 0.004 0.03 $\pm$ 0.003 0.03 $\pm$ 0.004 0.03 $\pm$ 0.006 0.03 $\pm$ 0.006
SP-P - Prob. Avg. 0.04 $\pm$ 0.002 0.04 $\pm$ 0.004 0.03 $\pm$ 0.002 0.03 $\pm$ 0.002 0.03 $\pm$ 0.004 0.03 $\pm$ 0.005
SVM DDPM - Log Prob. Avg. 0.11 $\pm$ 0.051 0.11 $\pm$ 0.053 0.11 $\pm$ 0.053 0.11 $\pm$ 0.053 0.11 $\pm$ 0.051 0.11 $\pm$ 0.051
DDPM - Prob. Avg. 0.11 $\pm$ 0.051 0.11 $\pm$ 0.053 0.11 $\pm$ 0.053 0.11 $\pm$ 0.053 0.11 $\pm$ 0.051 0.11 $\pm$ 0.051
SP-P - Log Prob. Avg. 0.04 $\pm$ 0.005 0.04 $\pm$ 0.007 0.03 $\pm$ 0.005 0.03 $\pm$ 0.003 0.03 $\pm$ 0.003 0.03 $\pm$ 0.003
SP-P - Prob. Avg. 0.04 $\pm$ 0.005 0.04 $\pm$ 0.006 0.03 $\pm$ 0.004 0.03 $\pm$ 0.003 0.03 $\pm$ 0.004 0.03 $\pm$ 0.003
Logistic Regression DDPM - Log Prob. Avg. 0.11 $\pm$ 0.056 0.10 $\pm$ 0.071 0.10 $\pm$ 0.071 0.09 $\pm$ 0.075 0.09 $\pm$ 0.078 0.09 $\pm$ 0.077
DDPM - Prob. Avg. 0.11 $\pm$ 0.056 0.10 $\pm$ 0.069 0.10 $\pm$ 0.069 0.10 $\pm$ 0.074 0.10 $\pm$ 0.077 0.10 $\pm$ 0.076
SP-P - Log Prob. Avg. 0.04 $\pm$ 0.009 0.03 $\pm$ 0.003 0.03 $\pm$ 0.004 0.02 $\pm$ 0.004 0.02 $\pm$ 0.003 0.02 $\pm$ 0.003
SP-P - Prob. Avg. 0.04 $\pm$ 0.009 0.03 $\pm$ 0.003 0.03 $\pm$ 0.003 0.03 $\pm$ 0.003 0.03 $\pm$ 0.004 0.02 $\pm$ 0.004
Table S10: Table of results from the German Credit dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator
1-NN DDPM - Log Prob. Avg. 0.41 $\pm$ 0.030 0.33 $\pm$ 0.027 0.35 $\pm$ 0.017 0.34 $\pm$ 0.021 0.32 $\pm$ 0.014 0.32 $\pm$ 0.018
DDPM - Prob. Avg. 0.41 $\pm$ 0.030 0.33 $\pm$ 0.027 0.29 $\pm$ 0.007 0.26 $\pm$ 0.010 0.25 $\pm$ 0.002 0.25 $\pm$ 0.005
SP-P - Log Prob. Avg. 0.37 $\pm$ 0.037 0.28 $\pm$ 0.033 0.27 $\pm$ 0.026 0.26 $\pm$ 0.024 0.24 $\pm$ 0.031 0.23 $\pm$ 0.028
SP-P - Prob. Avg. 0.37 $\pm$ 0.037 0.28 $\pm$ 0.033 0.23 $\pm$ 0.015 0.20 $\pm$ 0.011 0.19 $\pm$ 0.013 0.18 $\pm$ 0.013
5-NN DDPM - Log Prob. Avg. 0.24 $\pm$ 0.022 0.25 $\pm$ 0.013 0.24 $\pm$ 0.019 0.25 $\pm$ 0.025 0.25 $\pm$ 0.028 0.25 $\pm$ 0.029
DDPM - Prob. Avg. 0.24 $\pm$ 0.022 0.23 $\pm$ 0.011 0.22 $\pm$ 0.010 0.22 $\pm$ 0.012 0.22 $\pm$ 0.013 0.22 $\pm$ 0.012
SP-P - Log Prob. Avg. 0.23 $\pm$ 0.011 0.22 $\pm$ 0.019 0.21 $\pm$ 0.012 0.21 $\pm$ 0.010 0.20 $\pm$ 0.007 0.20 $\pm$ 0.014
SP-P - Prob. Avg. 0.23 $\pm$ 0.011 0.20 $\pm$ 0.016 0.19 $\pm$ 0.011 0.18 $\pm$ 0.010 0.18 $\pm$ 0.007 0.18 $\pm$ 0.009
Decision Tree DDPM - Log Prob. Avg. 0.44 $\pm$ 0.048 0.32 $\pm$ 0.045 0.31 $\pm$ 0.038 0.30 $\pm$ 0.030 0.29 $\pm$ 0.017 0.28 $\pm$ 0.032
DDPM - Prob. Avg. 0.44 $\pm$ 0.048 0.32 $\pm$ 0.045 0.26 $\pm$ 0.025 0.24 $\pm$ 0.015 0.23 $\pm$ 0.008 0.22 $\pm$ 0.012
SP-P - Log Prob. Avg. 0.33 $\pm$ 0.037 0.25 $\pm$ 0.018 0.25 $\pm$ 0.013 0.23 $\pm$ 0.024 0.21 $\pm$ 0.018 0.20 $\pm$ 0.018
SP-P - Prob. Avg. 0.33 $\pm$ 0.037 0.25 $\pm$ 0.018 0.20 $\pm$ 0.014 0.19 $\pm$ 0.016 0.17 $\pm$ 0.013 0.16 $\pm$ 0.011
Random Forest DDPM - Log Prob. Avg. 0.22 $\pm$ 0.014 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010
DDPM - Prob. Avg. 0.22 $\pm$ 0.014 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.009
SP-P - Log Prob. Avg. 0.19 $\pm$ 0.006 0.18 $\pm$ 0.011 0.17 $\pm$ 0.009 0.17 $\pm$ 0.008 0.17 $\pm$ 0.007 0.17 $\pm$ 0.008
SP-P - Prob. Avg. 0.19 $\pm$ 0.006 0.18 $\pm$ 0.011 0.17 $\pm$ 0.009 0.17 $\pm$ 0.008 0.17 $\pm$ 0.007 0.17 $\pm$ 0.007
MLP DDPM - Log Prob. Avg. 0.33 $\pm$ 0.024 0.30 $\pm$ 0.016 0.30 $\pm$ 0.020 0.29 $\pm$ 0.024 0.29 $\pm$ 0.022 0.28 $\pm$ 0.015
DDPM - Prob. Avg. 0.33 $\pm$ 0.024 0.28 $\pm$ 0.018 0.26 $\pm$ 0.013 0.25 $\pm$ 0.020 0.24 $\pm$ 0.019 0.24 $\pm$ 0.014
SP-P - Log Prob. Avg. 0.27 $\pm$ 0.037 0.24 $\pm$ 0.030 0.21 $\pm$ 0.019 0.19 $\pm$ 0.016 0.19 $\pm$ 0.018 0.18 $\pm$ 0.016
SP-P - Prob. Avg. 0.27 $\pm$ 0.037 0.21 $\pm$ 0.021 0.18 $\pm$ 0.009 0.17 $\pm$ 0.012 0.16 $\pm$ 0.013 0.16 $\pm$ 0.013
Gradient Boosting DDPM - Log Prob. Avg. 0.23 $\pm$ 0.021 0.22 $\pm$ 0.012 0.22 $\pm$ 0.009 0.22 $\pm$ 0.009 0.22 $\pm$ 0.008 0.21 $\pm$ 0.012
DDPM - Prob. Avg. 0.23 $\pm$ 0.021 0.22 $\pm$ 0.013 0.22 $\pm$ 0.008 0.22 $\pm$ 0.008 0.21 $\pm$ 0.007 0.21 $\pm$ 0.010
SP-P - Log Prob. Avg. 0.20 $\pm$ 0.018 0.18 $\pm$ 0.017 0.17 $\pm$ 0.012 0.17 $\pm$ 0.014 0.17 $\pm$ 0.011 0.16 $\pm$ 0.012
SP-P - Prob. Avg. 0.20 $\pm$ 0.018 0.18 $\pm$ 0.017 0.17 $\pm$ 0.012 0.17 $\pm$ 0.012 0.17 $\pm$ 0.010 0.16 $\pm$ 0.010
SVM DDPM - Log Prob. Avg. 0.21 $\pm$ 0.010 0.21 $\pm$ 0.009 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010
DDPM - Prob. Avg. 0.21 $\pm$ 0.010 0.21 $\pm$ 0.009 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010 0.21 $\pm$ 0.010
SP-P - Log Prob. Avg. 0.20 $\pm$ 0.008 0.18 $\pm$ 0.013 0.17 $\pm$ 0.013 0.17 $\pm$ 0.011 0.17 $\pm$ 0.008 0.17 $\pm$ 0.009
SP-P - Prob. Avg. 0.20 $\pm$ 0.008 0.18 $\pm$ 0.014 0.17 $\pm$ 0.012 0.17 $\pm$ 0.010 0.17 $\pm$ 0.008 0.17 $\pm$ 0.008
Logistic Regression DDPM - Log Prob. Avg. 0.23 $\pm$ 0.017 0.22 $\pm$ 0.008 0.21 $\pm$ 0.009 0.21 $\pm$ 0.014 0.22 $\pm$ 0.013 0.21 $\pm$ 0.013
DDPM - Prob. Avg. 0.23 $\pm$ 0.017 0.22 $\pm$ 0.006 0.21 $\pm$ 0.007 0.21 $\pm$ 0.011 0.21 $\pm$ 0.010 0.21 $\pm$ 0.011
SP-P - Log Prob. Avg. 0.20 $\pm$ 0.003 0.18 $\pm$ 0.011 0.17 $\pm$ 0.013 0.17 $\pm$ 0.011 0.17 $\pm$ 0.009 0.16 $\pm$ 0.010
SP-P - Prob. Avg. 0.20 $\pm$ 0.003 0.18 $\pm$ 0.012 0.17 $\pm$ 0.013 0.17 $\pm$ 0.010 0.17 $\pm$ 0.008 0.16 $\pm$ 0.008
Table S11: Table of results from predicting MSE on the ACS 2018 dataset. The numbers are the mean measured and predicted MSEs $\pm$ standard deviations from 3 repeats. The nans for linear regression are due to excluding some repeats that had an extremely large MSE ($\geq 10^{6}$).
m 1 2 4 8 16 32
Downstream Generator Predicted / Measured
Linear Regression DDPM Predicted 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.008 0.88 $\pm$ 0.006 0.88 $\pm$ 0.007 0.88 $\pm$ 0.008
Measured 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.005 0.87 $\pm$ 0.001 0.87 $\pm$ 0.003 0.87 $\pm$ 0.003
SP-P Predicted 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.006 0.78 $\pm$ nan nan nan
Measured 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.007 0.78 $\pm$ nan nan nan
Ridge Regression DDPM Predicted 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.008 0.87 $\pm$ 0.009 0.87 $\pm$ 0.010 0.87 $\pm$ 0.010
Measured 0.87 $\pm$ 0.008 0.87 $\pm$ 0.007 0.87 $\pm$ 0.005 0.87 $\pm$ 0.003 0.87 $\pm$ 0.004 0.87 $\pm$ 0.003
SP-P Predicted 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.006 0.77 $\pm$ 0.006 0.77 $\pm$ 0.006 0.77 $\pm$ 0.006
Measured 0.78 $\pm$ 0.007 0.78 $\pm$ 0.006 0.78 $\pm$ 0.007 0.78 $\pm$ 0.007 0.77 $\pm$ 0.007 0.77 $\pm$ 0.006
1-NN DDPM Predicted 1.55 $\pm$ 0.022 1.26 $\pm$ 0.026 1.11 $\pm$ 0.027 1.04 $\pm$ 0.028 1.01 $\pm$ 0.029 0.99 $\pm$ 0.029
Measured 1.55 $\pm$ 0.022 1.26 $\pm$ 0.026 1.10 $\pm$ 0.019 1.03 $\pm$ 0.014 1.00 $\pm$ 0.013 0.98 $\pm$ 0.011
SP-P Predicted 1.38 $\pm$ 0.015 1.04 $\pm$ 0.010 0.87 $\pm$ 0.009 0.79 $\pm$ 0.010 0.75 $\pm$ 0.010 0.72 $\pm$ 0.010
Measured 1.38 $\pm$ 0.015 1.04 $\pm$ 0.010 0.88 $\pm$ 0.001 0.79 $\pm$ 0.004 0.75 $\pm$ 0.006 0.73 $\pm$ 0.007
5-NN DDPM Predicted 0.98 $\pm$ 0.006 0.93 $\pm$ 0.009 0.91 $\pm$ 0.010 0.90 $\pm$ 0.011 0.89 $\pm$ 0.011 0.89 $\pm$ 0.012
Measured 0.98 $\pm$ 0.006 0.93 $\pm$ 0.009 0.90 $\pm$ 0.008 0.89 $\pm$ 0.007 0.89 $\pm$ 0.004 0.88 $\pm$ 0.004
SP-P Predicted 0.93 $\pm$ 0.002 0.82 $\pm$ 0.005 0.77 $\pm$ 0.008 0.74 $\pm$ 0.009 0.73 $\pm$ 0.009 0.72 $\pm$ 0.010
Measured 0.93 $\pm$ 0.002 0.82 $\pm$ 0.005 0.78 $\pm$ 0.004 0.75 $\pm$ 0.008 0.74 $\pm$ 0.008 0.73 $\pm$ 0.009
Decision Tree DDPM Predicted 1.41 $\pm$ 0.008 1.09 $\pm$ 0.005 0.94 $\pm$ 0.008 0.86 $\pm$ 0.009 0.82 $\pm$ 0.010 0.80 $\pm$ 0.011
Measured 1.41 $\pm$ 0.008 1.09 $\pm$ 0.005 0.95 $\pm$ 0.006 0.87 $\pm$ 0.004 0.84 $\pm$ 0.006 0.82 $\pm$ 0.011
SP-P Predicted 1.25 $\pm$ 0.035 0.96 $\pm$ 0.023 0.82 $\pm$ 0.018 0.75 $\pm$ 0.017 0.72 $\pm$ 0.016 0.70 $\pm$ 0.016
Measured 1.25 $\pm$ 0.035 0.96 $\pm$ 0.023 0.82 $\pm$ 0.011 0.74 $\pm$ 0.005 0.71 $\pm$ 0.003 0.69 $\pm$ 0.004
Random Forest DDPM Predicted 0.88 $\pm$ 0.010 0.82 $\pm$ 0.013 0.80 $\pm$ 0.015 0.78 $\pm$ 0.016 0.77 $\pm$ 0.017 0.77 $\pm$ 0.017
Measured 0.88 $\pm$ 0.010 0.82 $\pm$ 0.013 0.80 $\pm$ 0.009 0.78 $\pm$ 0.009 0.78 $\pm$ 0.009 0.77 $\pm$ 0.009
SP-P Predicted 0.84 $\pm$ 0.010 0.76 $\pm$ 0.010 0.72 $\pm$ 0.010 0.70 $\pm$ 0.010 0.69 $\pm$ 0.011 0.68 $\pm$ 0.011
Measured 0.84 $\pm$ 0.010 0.76 $\pm$ 0.010 0.72 $\pm$ 0.006 0.69 $\pm$ 0.005 0.68 $\pm$ 0.005 0.68 $\pm$ 0.006
Gradient Boosting DDPM Predicted 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008
Measured 0.74 $\pm$ 0.008 0.74 $\pm$ 0.008 0.74 $\pm$ 0.006 0.74 $\pm$ 0.005 0.74 $\pm$ 0.004 0.74 $\pm$ 0.004
SP-P Predicted 0.68 $\pm$ 0.005 0.68 $\pm$ 0.005 0.68 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.005
Measured 0.68 $\pm$ 0.005 0.68 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.005 0.67 $\pm$ 0.006 0.67 $\pm$ 0.006
MLP DDPM Predicted 0.82 $\pm$ 0.019 0.77 $\pm$ 0.018 0.75 $\pm$ 0.018 0.74 $\pm$ 0.019 0.74 $\pm$ 0.019 0.73 $\pm$ 0.019
Measured 0.82 $\pm$ 0.019 0.77 $\pm$ 0.018 0.76 $\pm$ 0.017 0.75 $\pm$ 0.007 0.74 $\pm$ 0.005 0.74 $\pm$ 0.004
SP-P Predicted 0.76 $\pm$ 0.006 0.71 $\pm$ 0.003 0.68 $\pm$ 0.003 0.67 $\pm$ 0.003 0.66 $\pm$ 0.004 0.66 $\pm$ 0.004
Measured 0.76 $\pm$ 0.006 0.71 $\pm$ 0.003 0.68 $\pm$ 0.008 0.67 $\pm$ 0.007 0.67 $\pm$ 0.007 0.66 $\pm$ 0.006
SVM DDPM Predicted 0.79 $\pm$ 0.004 0.78 $\pm$ 0.002 0.77 $\pm$ 0.004 0.77 $\pm$ 0.005 0.77 $\pm$ 0.005 0.77 $\pm$ 0.006
Measured 0.79 $\pm$ 0.004 0.78 $\pm$ 0.002 0.77 $\pm$ 0.002 0.78 $\pm$ 0.002 0.78 $\pm$ 0.003 0.78 $\pm$ 0.002
SP-P Predicted 0.71 $\pm$ 0.005 0.69 $\pm$ 0.003 0.69 $\pm$ 0.002 0.68 $\pm$ 0.001 0.68 $\pm$ 0.001 0.68 $\pm$ 0.001
Measured 0.71 $\pm$ 0.005 0.69 $\pm$ 0.003 0.69 $\pm$ 0.002 0.69 $\pm$ 0.004 0.69 $\pm$ 0.004 0.68 $\pm$ 0.004
Table S12: Table of results from predicting Brier score on the German Credit dataset. The numbers are the mean measured and predicted Brier scores $\pm$ standard deviations from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator Predicted / Measured
1-NN DDPM Measured 0.41 $\pm$ 0.027 0.33 $\pm$ 0.024 0.32 $\pm$ 0.032 0.30 $\pm$ 0.044 0.29 $\pm$ 0.036 0.28 $\pm$ 0.038
Predicted 0.41 $\pm$ 0.027 0.33 $\pm$ 0.024 0.30 $\pm$ 0.022 0.28 $\pm$ 0.021 0.27 $\pm$ 0.021 0.27 $\pm$ 0.021
SP-P Measured 0.37 $\pm$ 0.033 0.28 $\pm$ 0.029 0.25 $\pm$ 0.030 0.23 $\pm$ 0.035 0.21 $\pm$ 0.035 0.20 $\pm$ 0.033
Predicted 0.37 $\pm$ 0.033 0.28 $\pm$ 0.029 0.24 $\pm$ 0.037 0.22 $\pm$ 0.043 0.21 $\pm$ 0.046 0.20 $\pm$ 0.047
5-NN DDPM Measured 0.24 $\pm$ 0.020 0.24 $\pm$ 0.015 0.23 $\pm$ 0.019 0.23 $\pm$ 0.024 0.23 $\pm$ 0.025 0.23 $\pm$ 0.025
Predicted 0.24 $\pm$ 0.020 0.24 $\pm$ 0.015 0.23 $\pm$ 0.017 0.23 $\pm$ 0.020 0.23 $\pm$ 0.021 0.23 $\pm$ 0.022
SP-P Measured 0.23 $\pm$ 0.009 0.21 $\pm$ 0.018 0.20 $\pm$ 0.017 0.19 $\pm$ 0.018 0.19 $\pm$ 0.015 0.19 $\pm$ 0.018
Predicted 0.23 $\pm$ 0.009 0.21 $\pm$ 0.018 0.20 $\pm$ 0.024 0.20 $\pm$ 0.026 0.19 $\pm$ 0.028 0.19 $\pm$ 0.029
Decision Tree DDPM Measured 0.44 $\pm$ 0.043 0.32 $\pm$ 0.040 0.28 $\pm$ 0.038 0.27 $\pm$ 0.040 0.26 $\pm$ 0.037 0.25 $\pm$ 0.040
Predicted 0.44 $\pm$ 0.043 0.32 $\pm$ 0.040 0.26 $\pm$ 0.039 0.23 $\pm$ 0.039 0.21 $\pm$ 0.038 0.20 $\pm$ 0.038
SP-P Measured 0.33 $\pm$ 0.033 0.25 $\pm$ 0.016 0.23 $\pm$ 0.027 0.21 $\pm$ 0.031 0.19 $\pm$ 0.026 0.18 $\pm$ 0.022
Predicted 0.33 $\pm$ 0.033 0.25 $\pm$ 0.016 0.22 $\pm$ 0.007 0.20 $\pm$ 0.003 0.19 $\pm$ 0.001 0.18 $\pm$ 0.001
Random Forest DDPM Measured 0.22 $\pm$ 0.013 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009
Predicted 0.22 $\pm$ 0.013 0.21 $\pm$ 0.008 0.21 $\pm$ 0.006 0.21 $\pm$ 0.005 0.21 $\pm$ 0.004 0.21 $\pm$ 0.004
SP-P Measured 0.19 $\pm$ 0.005 0.18 $\pm$ 0.010 0.17 $\pm$ 0.008 0.17 $\pm$ 0.007 0.17 $\pm$ 0.006 0.17 $\pm$ 0.007
Predicted 0.19 $\pm$ 0.005 0.18 $\pm$ 0.010 0.17 $\pm$ 0.012 0.17 $\pm$ 0.014 0.17 $\pm$ 0.014 0.17 $\pm$ 0.015
MLP DDPM Measured 0.33 $\pm$ 0.022 0.29 $\pm$ 0.019 0.28 $\pm$ 0.025 0.27 $\pm$ 0.032 0.26 $\pm$ 0.030 0.26 $\pm$ 0.026
Predicted 0.33 $\pm$ 0.022 0.29 $\pm$ 0.019 0.28 $\pm$ 0.023 0.27 $\pm$ 0.026 0.26 $\pm$ 0.027 0.26 $\pm$ 0.028
SP-P Measured 0.27 $\pm$ 0.033 0.23 $\pm$ 0.027 0.19 $\pm$ 0.022 0.18 $\pm$ 0.019 0.18 $\pm$ 0.020 0.17 $\pm$ 0.018
Predicted 0.27 $\pm$ 0.033 0.23 $\pm$ 0.027 0.21 $\pm$ 0.044 0.20 $\pm$ 0.054 0.19 $\pm$ 0.059 0.19 $\pm$ 0.061
Gradient Boosting DDPM Measured 0.23 $\pm$ 0.019 0.22 $\pm$ 0.011 0.22 $\pm$ 0.008 0.22 $\pm$ 0.008 0.22 $\pm$ 0.007 0.21 $\pm$ 0.010
Predicted 0.23 $\pm$ 0.019 0.22 $\pm$ 0.011 0.22 $\pm$ 0.008 0.22 $\pm$ 0.007 0.22 $\pm$ 0.007 0.22 $\pm$ 0.007
SP-P Measured 0.20 $\pm$ 0.016 0.18 $\pm$ 0.016 0.17 $\pm$ 0.011 0.17 $\pm$ 0.012 0.17 $\pm$ 0.009 0.16 $\pm$ 0.010
Predicted 0.20 $\pm$ 0.016 0.18 $\pm$ 0.016 0.17 $\pm$ 0.018 0.17 $\pm$ 0.020 0.16 $\pm$ 0.021 0.16 $\pm$ 0.022
SVM DDPM Measured 0.21 $\pm$ 0.009 0.21 $\pm$ 0.008 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009
Predicted 0.21 $\pm$ 0.009 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008
SP-P Measured 0.20 $\pm$ 0.007 0.18 $\pm$ 0.012 0.17 $\pm$ 0.011 0.17 $\pm$ 0.010 0.17 $\pm$ 0.007 0.17 $\pm$ 0.008
Predicted 0.20 $\pm$ 0.007 0.18 $\pm$ 0.012 0.17 $\pm$ 0.016 0.17 $\pm$ 0.018 0.17 $\pm$ 0.019 0.17 $\pm$ 0.020
Logistic Regression DDPM Measured 0.23 $\pm$ 0.015 0.22 $\pm$ 0.006 0.21 $\pm$ 0.007 0.21 $\pm$ 0.011 0.21 $\pm$ 0.011 0.21 $\pm$ 0.011
Predicted 0.23 $\pm$ 0.015 0.22 $\pm$ 0.006 0.21 $\pm$ 0.003 0.21 $\pm$ 0.003 0.21 $\pm$ 0.004 0.21 $\pm$ 0.004
SP-P Measured 0.20 $\pm$ 0.003 0.18 $\pm$ 0.010 0.17 $\pm$ 0.012 0.17 $\pm$ 0.009 0.17 $\pm$ 0.008 0.16 $\pm$ 0.008
Predicted 0.20 $\pm$ 0.003 0.18 $\pm$ 0.010 0.17 $\pm$ 0.017 0.17 $\pm$ 0.020 0.16 $\pm$ 0.021 0.16 $\pm$ 0.022
Table S13: Table of results from the DP experiment. The numbers are the mean Brier score $\pm$ the standard deviation from 3 repeats. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$.
m 1 2 4 8 16 32
Downstream Generator
1-NN AIM - Log Prob. Avg. 0.23 $\pm$ 0.012 0.17 $\pm$ 0.003 0.16 $\pm$ 0.002 0.16 $\pm$ 0.011 0.15 $\pm$ 0.006 0.15 $\pm$ 0.001
AIM - Prob. Avg. 0.23 $\pm$ 0.012 0.17 $\pm$ 0.003 0.14 $\pm$ 0.003 0.13 $\pm$ 0.004 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.25 $\pm$ 0.008 0.20 $\pm$ 0.020 0.19 $\pm$ 0.005 0.17 $\pm$ 0.006 0.16 $\pm$ 0.005 0.16 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.25 $\pm$ 0.008 0.20 $\pm$ 0.020 0.16 $\pm$ 0.003 0.14 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003
5-NN AIM - Log Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.002 0.14 $\pm$ 0.002 0.14 $\pm$ 0.002 0.13 $\pm$ 0.004 0.13 $\pm$ 0.001
AIM - Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.16 $\pm$ 0.007 0.15 $\pm$ 0.004 0.15 $\pm$ 0.008 0.15 $\pm$ 0.008 0.14 $\pm$ 0.003 0.14 $\pm$ 0.003
NAPSU-MQ - Prob. Avg. 0.16 $\pm$ 0.007 0.14 $\pm$ 0.005 0.13 $\pm$ 0.005 0.13 $\pm$ 0.005 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
Decision Tree AIM - Log Prob. Avg. 0.13 $\pm$ 0.003 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.16 $\pm$ 0.005 0.15 $\pm$ 0.004 0.15 $\pm$ 0.005 0.14 $\pm$ 0.004 0.14 $\pm$ 0.004 0.14 $\pm$ 0.003
NAPSU-MQ - Prob. Avg. 0.16 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
Random Forest AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.007 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.007 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003
MLP AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.006 0.13 $\pm$ 0.004 0.13 $\pm$ 0.005 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.006 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.12 $\pm$ 0.004
Gradient Boosting AIM - Log Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004
SVM AIM - Log Prob. Avg. 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.001 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002
AIM - Prob. Avg. 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.001 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.005 0.14 $\pm$ 0.006 0.14 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.005 0.13 $\pm$ 0.005
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.005 0.14 $\pm$ 0.006 0.14 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.005
Logistic Regression AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.13 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.13 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
E.4 Estimating Model and Synthetic Data Variances
In this section, we estimate the MV and SDV terms from the decomposition in Theorem 2.2. We first generate 32 synthetic datasets that are 5 times larger than the real dataset, and split each synthetic datasets into 5 equally-sized subsets. This is equivalent to sampling 32 $\theta_{i}$ values, and for each $\theta_{i}$, generating 5 synthetic datasets i.i.d.  We then train the downstream predictor on each synthetic dataset, and store the predictions for all test points.
To estimate MV, we compute the sample variance over the 5 synthetic datasets generated from the same $\theta_{i}$, and then compute the mean over the 32 different $\theta_{i}$ values. To estimate SDV, we compute the sample mean over the 5 synthetic datasets from the same $\theta_{i}$, and compute the sample variance over the 32 different $\theta_{i}$ values.
The result is an estimate of MV and SDV for each test point. We plot the mean over the test points. The whole experiment is repeated 3 times, with different train-test splits. The datasets, train-test splits, and downstream predictors are the same as in the other experiments, described in Appendix D.
The results are in Figure S11. MV depends mostly on the downstream predictor, while SDV also depends on the synthetic data generator. We also confirm that decision trees and 1-NN have much higher variance than the other models, and linear, ridge and logistic regression have a very low variance.
(a) Abalone
(b) ACS 2018
(c) Adult
(d) Breast Cancer
(e) California Housing
(f) German Credit
(g) Insurance
Figure S11: Estimating the MV and SDV terms from the decomposition. Decision trees have high variances on all datasets, while linear, ridge and logistic regression have low variances. MV depends mostly on the predictor, while SDV depends on both the predictor and synthetic data generation algorithm. The points are the averages of estimated MV and SDV, averaged over the test data, from 3 repeats with different train-test splits. We excluded some test data points that had extremely large variance estimates ($\geq 10^{6}$) for linear regression on the ACS 2018 dataset. The predictors are decision tree (DT), nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), linear regression (LR), ridge regression (RR) and logistic regression (LogR). The synthetic data generators are DDPM and synthpop (SP-P).